### v1

# Optimal Control of a Subdiffusion model using Dirichlet-Neumann and Neumann-Neumann Waveform Relaxation Algorithms

Abstract
This paper investigates the convergence behavior of the Dirichlet-Neumann and Neumann-Neumann waveform relaxation algorithms for optimal control problem with sub-diffusion PDE as constraint. The algorithms are applied to regular domains in 1D for multiple subdomains, and the impact of different constant values of the generalized diffusion coefficient on the algorithms’ convergence is analyzed. The convergence rate of the algorithms is analyzed as the fractional order of the time derivative changes.
1 Optimal Control of a subdiffusion model
Let, $\Omega$ be the open bounded subset in $\mathbb{R}^{d}$ with boundary $\partial\Omega$, and $[0,T]$ is the time domain with $T>0$. First consider the quadratic cost functional
(1) $J(y,u):=\frac{1}{2}\|y-y_{Q}\|^{2}_{L^{2}((0,T)\times\Omega)}+\frac{\sigma}{2}\|u\|^{2}_{L^{2}((0,T)\times\Omega)},$
where $u(\boldsymbol{x},t)$ is the control variable and $y_{Q}(\boldsymbol{x},t)$ be the target state. $\sigma$ be the regularization parameter, which mainly use to penalize the control variable. Our objective is to achieve the target state $y_{Q}(\boldsymbol{x},t)$ using the control $u(\boldsymbol{x},t)$, so that the cost functional $J(y,u)$ will be minimum. Here, the state variable $y(\boldsymbol{x},t)$ is governed by the fractional diffusion equation of order $\alpha>0$ given as follows:
(2) $\begin{cases}D^{\alpha}_{RL}y=\nabla\cdot\left(\kappa(\boldsymbol{x},t)\nabla y\right)+f(\boldsymbol{x},t)+u,&\textrm{in}\;\Omega\times(0,T),\\ y(\boldsymbol{x},t)=g(\boldsymbol{x},t),&\textrm{on}\;\partial\Omega\times(0,T),\\ I^{1-\alpha}_{RL}y(\boldsymbol{x},0^{+})=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega.\end{cases}$
Here, $\kappa(\boldsymbol{x},t)>0$ is the dimensionless diffusion coefficient and $D^{\alpha}_{RL}$ be the left sided Riemann-Liouville fractional derivative [7] defined for order $\alpha$, $m-1<\alpha<m$ and $m\in\mathbb{N}$ as follows:
[…]
And $I^{\alpha}_{RL}$ represents the Riemann-Liouville fractional integral of order $\alpha>0$ on a continuous function $f(t)$ defined as follows:
[…]
For $10<\alpha<$ 1 and $f,u\in L^{2}((0,T)\times\Omega)$ the problem (2) has a unique solution $u\in L^{2}((0,T);H^{2}(\Omega)\cap H^{1}_{0}(\Omega))$ for zero Dirichlet boundary condition, see  [5]. And $J$ be the convex function, so unique infimum exist. If $u\in L^{2}((0,T)\times\Omega)$ be the solution of (1) then using the Euler-Lagrange optimality conditions we have adjoint variable $p\in L^{2}((0,T)\times\Omega)$ such that $(u,y,p)$ satisfy the optimality system  [4]:
(3) $\begin{cases}D^{\alpha}_{RL}y=\nabla\cdot\left(\kappa(\boldsymbol{x},t)\nabla y\right)+f(\boldsymbol{x},t)+u,&\textrm{in}\;\Omega\times(0,T),\\ y(\boldsymbol{x},t)=g(\boldsymbol{x},t),&\textrm{on}\;\partial\Omega\times(0,T),\\ I^{1-\alpha}_{+}y(\boldsymbol{x},0)=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega.\end{cases}$
(4) $\begin{cases}{}_{T}D^{\alpha}_{t}p=\nabla\cdot\left(\kappa(\boldsymbol{x},t)\nabla p\right)+y-y_{Q},&\textrm{in}\;\Omega\times(0,T),\\ p(\boldsymbol{x},t)=0,&\textrm{on}\;\partial\Omega\times(0,T),\\ p(\boldsymbol{x},T)=0,&\textrm{in}\;\Omega.\end{cases}$
with the condition $\sigma u=-p$. Here ${}_{T}D^{\alpha}_{t}$ be the right sided Caputo fractional derivative [7] defined for order $\alpha$, $m-1<\alpha<m$ and $m\in\mathbb{N}$ as follows:
[…]
2 The Dirichlet-Neumann Waveform Relaxation algorithm
The DNWR method [2] is a semi-parallel iterative algorithm, with the exception of the two sub-domain case. This algorithm combines the substructuring DD method for space and the waveform relaxation method for time. Let us consider two sub-domains $\Omega_{1}$ and $\Omega_{2}$ of $\Omega$ that do not overlap, with the interface $\Gamma:=\partial\Omega_{1}\cap\partial\Omega_{2}$. The solution $u$ is restricted to $u_{j}$ on $\Gamma$, where $j=1,2$, and $\boldsymbol{n}_{j}$ represents the unit outward normal on $\Gamma$ for $\Omega{j}$.
An initial guess $\xi^{(0)}(\boldsymbol{x},t)$ for state equation and $\zeta^{(0)}(\boldsymbol{x},t)$ for adjoint equation are selected throughout the interface $\Gamma\times(0,T)$. The solutions $(y_{j}^{(k)},p_{j}^{(k)}),j=1,2,\text{ for }k=1,2,\ldots$ are then obtained through the following Dirichlet-Neumann steps.
(5) $\begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}y_{1}^{(k)}=\nabla\cdot\left(\kappa_{1}\nabla y_{1}^{(k)}\right)+f-\frac{p_{1}^{(k)}}{\sigma},&\textrm{in}\;\Omega_{1},\\ I^{1-\alpha}_{RL}y_{1}^{(k)}(\boldsymbol{x},0)=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega_{1},\\ y_{1}^{(k)}=\xi^{(k-1)},&\textrm{on}\;\Gamma,\\ y_{1}^{(k)}=g,&\textrm{on}\;\partial\Omega_{1}\setminus\Gamma,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}y_{2}^{(k)}=\nabla\cdot\left(\kappa_{2}\nabla y_{2}^{(k)}\right)+f-\frac{p_{2}^{(k)}}{\sigma},&\textrm{in}\;\Omega_{2},\\ I^{1-\alpha}_{RL}y_{2}^{(k)}(\boldsymbol{x},0)=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega_{2},\\ \kappa_{2}\partial_{\boldsymbol{n}_{2}}y_{2}^{(k)}=-\kappa_{1}\partial_{\boldsymbol{n}_{1}}y_{1}^{(k)},&\textrm{on}\;\Gamma,\\ y_{2}^{(k)}=g,&\textrm{on}\;\partial\Omega_{2}\setminus\Gamma,\end{cases}\end{array}$
and the adjoint problem will be:
(6) $\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{1}^{(k)}=\nabla\cdot\left(\kappa_{1}\nabla p_{1}^{(k)}\right)+y_{1}^{(k)}-y_{Q}|_{1},&\textrm{in}\;\Omega_{1},\\ p_{1}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;\Omega_{1},\\ p_{1}^{(k)}=\zeta^{(k-1)},&\textrm{on}\;\Gamma,\\ p_{1}^{(k)}=g,&\textrm{on}\;\partial\Omega_{1}\setminus\Gamma,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{2}^{(k)}=\nabla\cdot\left(\kappa_{2}\nabla p_{2}^{(k)}\right)+y_{2}^{(k)}-y_{Q}|_{2},&\textrm{in}\;\Omega_{2},\\ p_{2}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;\Omega_{2},\\ \kappa_{2}\partial_{\boldsymbol{n}_{2}}p_{2}^{(k)}=-\kappa_{1}\partial_{\boldsymbol{n}_{1}}p_{1}^{(k)},&\textrm{on}\;\Gamma,\\ p_{2}^{(k)}=g,&\textrm{on}\;\partial\Omega_{2}\setminus\Gamma,\end{cases}\end{array}$
and then update the interface with the relaxation parameter $\theta\in(0,1]$ by
(7) $\xi^{(k)}(\boldsymbol{x},t)$ $=\theta u_{2}^{(k)}\left|{}_{\Gamma\times(0,T)}\right.+(1-\theta)\xi^{(k-1)}(\boldsymbol{x},t),$
(8) $\zeta^{(k)}(\boldsymbol{x},t)$ $=\phi p_{2}^{(k)}\left|{}_{\Gamma\times(0,T)}\right.+(1-\phi)\zeta^{(k-1)}(\boldsymbol{x},t),$
The primary objective of our analysis is to investigate how the update component error $\omega^{(k-1)}(\boldsymbol{x},t):=y|_{\Gamma\times(0,T)}-\xi^{(k-1)}(\boldsymbol{x},t)$, and $\upsilon^{(k-1)}(\boldsymbol{x},t):=p|_{\Gamma\times(0,T)}-\zeta^{(k-1)}(\boldsymbol{x},t)$, where $u|_{\Gamma\times(0,T)}$ and $p|_{\Gamma\times(0,T)}$ are the exact solution on the interface, converges to zero. Linearity suffices to consider $f(\boldsymbol{x},t)=0$, $g(\boldsymbol{x},t)=0$ for the convergence of error equation with $u_{0}(\boldsymbol{x})=0$ and $\upsilon_{0}(\boldsymbol{x})=0$ in (5) and (6) respectively. This will be studied in section 6.
3 The Neumann-Neumann Waveform Relaxation algorithm
We are considering the NNWR for multiple subdomains setting, where the model problem (3) and (4) on spatial domain $\Omega$ is divided into multiple non-overlapping subdomains $\Omega_{i}$, where $1\leq i\leq N$, and without cross points. We define $\Gamma_{i,j}$ as the intersection of $\partial\Omega_{i}$ and $\partial\Omega_{j}$ for $j=i+1,i-1$. On each subdomain $\Omega_{i}$, we employ the use of unit outward normals $\boldsymbol{n}{i,i-1}$ and $\boldsymbol{n}{i,i+1}$ on the interfaces $\Gamma_{i,i-1}$ and $\Gamma_{i,i+1}$ respectively. Given the initial guess $\xi_{i,j}^{(0)}(\boldsymbol{x},t)$ for state equation and $\zeta_{i,j}^{(0)}(\boldsymbol{x},t)$ for adjoint equation, along the interfaces $\Gamma_{i,j}\times(0,T)$, the NNWR algorithm can be expressed as follows:
(9) $\begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}y_{i}^{(k)}=\nabla\cdot\left(\kappa_{i}\nabla y_{i}^{(k)}\right)+f-\frac{p_{i}^{(k)}}{\sigma},&\mbox{in $\Omega_{i}$},\\ I^{1-\alpha}_{RL}y_{i}^{(k)}(\boldsymbol{x},0)=y_{0}(\boldsymbol{x}),&\mbox{in$\Omega_{i}$},\\ y_{i}^{(k)}=g,&\mbox{on $\partial\Omega_{i}\cap\partial\Omega$},\\ y_{i}^{(k)}=\xi_{i,j}^{(k-1)},&\mbox{on $\Gamma_{i,j}$},\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}\psi_{i}^{(k)}=\nabla\cdot\left(\kappa_{i}\nabla\psi_{i}^{(k)}\right)-\frac{\chi_{i}^{(k)}}{\sigma},&\mbox{in $\Omega_{i}$},\\ I^{1-\alpha}_{RL}\psi_{i}^{(k)}(\boldsymbol{x},0)=0,&\mbox{in $\Omega_{i}$},\\ \psi_{i}^{(k)}=0,&\mbox{on $\partial\Omega_{i}\cap\partial\Omega$},\\ \kappa_{i}\partial_{\boldsymbol{n}_{i,j}}\psi_{i}^{(k)}=\kappa_{i}\partial_{\boldsymbol{n}_{i,j}}y_{i}^{(k)}+\kappa_{j}\partial_{\boldsymbol{n}_{j,i}}y_{j}^{(k)},&\mbox{on $\Gamma_{i,j}$}.\end{cases}\end{array}$
(10) $\begin{array}[]{rcll}\begin{cases}{}_{T}D^{\alpha}_{t}p_{i}^{(k)}=\nabla\cdot\left(\kappa_{i}\nabla p_{i}^{(k)}\right)+y_{i}^{(k)}-y_{Q}|_{i},&\mbox{in $\Omega_{i}$},\\ p_{i}^{(k)}(\boldsymbol{x},T)=0,&\mbox{in $\Omega_{i}$},\\ p_{i}^{(k)}=0,&\mbox{on $\partial\Omega_{i}\cap\partial\Omega$},\\ p_{i}^{(k)}=\zeta_{i,j}^{(k-1)},&\mbox{on $\Gamma_{i,j}$},\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}{}_{T}D^{\alpha}_{t}\chi_{i}^{(k)}=\nabla\cdot\left(\kappa_{i}\nabla\chi_{i}^{(k)}\right)+\psi_{i}^{(k)},&\mbox{in $\Omega_{i}$},\\ \chi_{i}^{(k)}(\boldsymbol{x},T)=0,&\mbox{in $\Omega_{i}$},\\ \chi_{i}^{(k)}=0,&\mbox{on $\partial\Omega_{i}\cap\partial\Omega$},\\ \kappa_{i}\partial_{\boldsymbol{n}_{i,j}}\chi_{i}^{(k)}=\kappa_{i}\partial_{\boldsymbol{n}_{i,j}}p_{i}^{(k)}+\kappa_{j}\partial_{\boldsymbol{n}_{j,i}}p_{j}^{(k)},&\mbox{on $\Gamma_{i,j}$}.\end{cases}\end{array}$
The formula is then used to update the interface values:
(11) $\xi_{i,j}^{(k)}(\boldsymbol{x},t)$ $=\xi_{i,j}^{(k-1)}(\boldsymbol{x},t)-\theta_{i,j}\left(\psi_{i}^{(k)}\left|{}_{\Gamma_{i,j}\times(0,T)}\right.+\psi_{j}^{(k)}\left|{}_{\Gamma_{i,j}\times(0,T)}\right.\right),$
(12) $\zeta_{i,j}^{(k)}(\boldsymbol{x},t)$ $=\zeta_{i,j}^{(k-1)}(\boldsymbol{x},t)-\phi_{i,j}\left(\chi_{i}^{(k)}\left|{}_{\Gamma_{i,j}\times(0,T)}\right.+\chi_{j}^{(k)}\left|{}_{\Gamma_{i,j}\times(0,T)}\right.\right),$
Our goal is to study the convergence of error terms on artificial boundary $\omega_{i,j}^{(k-1)}(\boldsymbol{x},t):=y_{i}|_{\Gamma_{i,j}\times(0,T)}-\xi_{i,j}^{(k-1)}(\boldsymbol{x},t)$ and $\upsilon_{i,j}^{(k-1)}(\boldsymbol{x},t):=p_{i}|_{\Gamma_{i,j}\times(0,T)}-\zeta_{i,j}^{(k-1)}(\boldsymbol{x},t)$ as $k\to\infty$.
4 Auxiliary Results
Now we need some auxiliary result which are necessary for the followings convergent result of DNWR and NNWR algorithm.
Lemma 4.1 (Theorem 2.10 [5]).
Riemann-Liouville time derivative $D^{\alpha}_{RL}y(t)$ with homogeneous initial condition $I^{1-\alpha}_{RL}y(0^{+})=0$ is equivalent to Caputo fractional time derivative $D^{\alpha}_{C}y(t)$ with zero initial condition $y(0)=0$.
Proof 4.2.
We know from the relation between Riemann-Liouville and Caputo fractional derivative that
$D^{\alpha}_{C}y(t)=D^{\alpha}_{RL}y(t)+\frac{t^{-\alpha}}{\Gamma(1-\alpha)}y(0).$
After integrating both side with order $\alpha$ i.e. $I^{\alpha}$ we have $y(t)-y(0)=y(t)-\frac{t^{-\alpha}}{\Gamma(1-\alpha)}I^{(1-\alpha)}y(0)$. Thus $I^{(1-\alpha)}y(0)=0$ implies $y(0)=0$. Inversely from the same relation $y(0)=0$ implies $D^{\alpha}_{C}y(t)=D^{\alpha}_{RL}y(t)$. Hence equivalent.
Lemma 4.3.
Stability of an elliptic pde depends on the absolute minimum eigenvalue of the discretize spatial matrix.
Proof 4.4.
For sake of simplicity consider the 2D Poisson equation $u_{xx}+u_{yy}=f(x,y)$. Any numerical scheme of this pde will be stable if the truncation error coming from the Taylor series expansion of the derivative does not effect the solution error too much. That is if we write in terms of error equation using the central difference scheme, we have:
$\frac{e_{i+1,j}-2e_{i,j}+e_{i-1,j}}{\Delta x^{2}}+\frac{e_{i,j+1}-2e_{i,j}+e_{i,j-1}}{\Delta y^{2}}=\eta_{i,j}.$
Therefore in matrix form we have $Ae=\eta$, where $e$ and $\eta$ are column vectors. Hence $e=A^{-1}\eta$. Therefore the stability depends on the absolute minimum eigenvalue of matrix $A$.
Lemma 4.5.
Eigenvalues of matrix $M=\begin{bmatrix}A&\frac{1}{\sigma}B\\ -B&A\end{bmatrix}$ are the collection of eigenvalues of $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$, where $A$ and $B$ are block matrix of order $n$.
Proof 4.6.
We know
$\begin{vmatrix}A&\frac{1}{\sigma}B\\ -B&A\end{vmatrix}=\frac{1}{\sqrt{\sigma^{n}}}\begin{vmatrix}\sqrt{\sigma}A&\frac{1}{\sqrt{\sigma}}B\\ -B&A\end{vmatrix}=\frac{1}{\sqrt{\sigma^{n}}}\begin{vmatrix}\sqrt{\sigma}A+iB&\frac{1}{\sqrt{\sigma}}B-iA\\ -B&A\end{vmatrix}$
Then using the determinant formula when $A$ is invertiable, we have
$\begin{vmatrix}A&B\\ C&D\end{vmatrix}=det(A)det(D-CA^{-1}B)$
we have
$\frac{1}{\sqrt{\sigma^{n}}}\begin{vmatrix}\sqrt{\sigma}L+iB&\frac{1}{\sqrt{\sigma}}B-iL\\ -B&L\end{vmatrix}$ $=\frac{1}{\sqrt{\sigma^{n}}}det\left(\sqrt{\sigma}L+iB\right)det\left(L+B(\sqrt{\sigma}L+iB)^{-1}\left(\frac{1}{\sqrt{\sigma}}B-iL\right)\right)$
$=det\left(L+\frac{i}{\sqrt{\sigma}}B\right)det\left(L-\frac{i}{\sqrt{\sigma}}B\right).$
Therefore, when $A$ is invertiable we have
$det\left(\begin{bmatrix}A&\frac{1}{\sigma}B\\ -B&A\end{bmatrix}\right)=det\left(A+\frac{i}{\sqrt{\sigma}}B\right)det\left(A-\frac{i}{\sqrt{\sigma}}B\right)$
But this formula does not contain any $A$ inverse, so we could extend it when $A$ is not invertiable. For this we need to consider the polynomial $P(x)$ of degree $2n$ defined by
$P(x):=det\left(\begin{bmatrix}A-xI_{n}&\frac{1}{\sigma}B\\ -B&A-xI_{n}\end{bmatrix}\right)-det\left(A-xI_{n}+\frac{i}{\sqrt{\sigma}}B\right)det\left(A-xI_{n}-\frac{i}{\sqrt{\sigma}}B\right)$
which has infinitely many zeros when $x$ is not the eigenvalue of $M$, therefore $P(x)$ is a zero polynomial. Hence
$det\left(\begin{bmatrix}A-xI_{n}&\frac{1}{\sigma}B\\ -B&A-xI_{n}\end{bmatrix}\right)=det\left(A-xI_{n}+\frac{i}{\sqrt{\sigma}}B\right)det\left(A-xI_{n}-\frac{i}{\sqrt{\sigma}}B\right).$
Which implies the eigenvalues of matrix $M$ are the collection of eigenvalues of $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$.
From the above result we know that the matrix $M$ has the same set of eigenvalues as of $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$. Now we want to know is there any connection among them or not. The following result gives the in depth understanding about that.
Lemma 4.7.
The matrix $M=\begin{bmatrix}A&\frac{1}{\sigma}B\\ -B&A\end{bmatrix}$ is similar to the matrix $\bar{M}=\begin{bmatrix}A+\frac{i}{\sqrt{\sigma}}B&0\\ 0&A-\frac{i}{\sqrt{\sigma}}B\end{bmatrix}$.
Proof 4.8.
Let $X=\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}$ be the block matrix of order $2n$. Now if there exist an invertible matrix $X$ such that $MX=X\bar{M}$ then we say that $M$ and $\bar{M}$ are similar. Therefore,
$\begin{bmatrix}A&\frac{1}{\sigma}B\\ -B&A\end{bmatrix}\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}$ $=\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}\begin{bmatrix}A+\frac{i}{\sqrt{\sigma}}B&0\\ 0&A-\frac{i}{\sqrt{\sigma}}B\end{bmatrix}$
$\begin{bmatrix}AX_{11}+\frac{1}{\sigma}BX_{21}&AX_{12}+\frac{1}{\sigma}BX_{22}\\ AX_{21}-BX_{11}&AX_{22}-BX_{12}\end{bmatrix}$ $=\begin{bmatrix}X_{11}(A+\frac{i}{\sqrt{\sigma}}B)&X_{12}(A-\frac{i}{\sqrt{\sigma}}B)\\ X_{21}(A+\frac{i}{\sqrt{\sigma}}B)&X_{22}(A-\frac{i}{\sqrt{\sigma}}B)\end{bmatrix}$
Now if we choose $X_{11}=I,X_{12}=I,X_{21}=i\sqrt{\sigma}I,X_{22}=-i\sqrt{\sigma}I$, where $I$ is the identity matrix of order $n$ then left and right side will be equal. Hence $X=\begin{bmatrix}I&I\\ i\sqrt{\sigma}I&-i\sqrt{\sigma}I\end{bmatrix}$ which is invertiable so the matrix $M$ is similar to matrix $\bar{M}$ by the matrix by $X$.
The above result conclude that the diagonalization of matrix $M$ is separately depends on the diagonalization of matrix block $A+\frac{i}{\sqrt{\sigma}}B$ and $A-\frac{i}{\sqrt{\sigma}}B$. The next result give the lower bound on the real part of eigenvalues.
Lemma 4.9.
Let $M=A+iB$ , where $A,B$ are square matrices of order $n$ in $\mathbb{R}$ and $B=B^{T}$. If the matrix $(A+A^{T})$ is positive definite then the real part of the eigenvalues of $M$ are strictly positive.
Proof 4.10.
To prove this result we use the general version of Lyapunov’s theorem on stability [6], which state that: Let $X$ be a complex matrix of order $n$ and $Y$ be a positive definite Hermitian matrix of order $n$. Then there exist a negative definite Hermitian matrix $Z$ for which $XZ+ZX^{*}=Y$ (${}^{\prime}*^{\prime}$ imply the conjugate transpose) holds, if and only if all the eigenvalues of $X$ have negative real part.
Now if we choose $Z=-I$ i.e. negative identity matrix and $X=-M$ then we have $Y=(A+iB)+(A^{T}-iB^{T})=A+A^{T}$ as $B=B^{T}$. From assumption $(A+A^{T})$ is positive definite so the real part of the eigenvalues of $-M$ are negative. Hence the real part of the eigenvalues of $M$ are strictly positive.
Corollary 4.11.
Let $A$ is the first order backward difference matrix with step size $\Delta t$ for $n$ nodes and $B$ is the unit anti-diagonal matrix of same order, i.e.
$A=\begin{bmatrix}1&&&&\\ -1&1&&&\\ &\ddots&\ddots&&\\ &&-1&1&\\ &&&-1&1\end{bmatrix},\hskip 28.45274pt\textit{ and }\hskip 28.45274ptB=\begin{bmatrix}&&&&1\\ &&&1&\\ &&\iddots&&\\ &1&&&\\ 1&&&&\end{bmatrix}.$
Clearly $B$ is symmetric and $-(A+A^{T})$ is the central difference matrix with the eigenvalues $\lambda_{j}=-\frac{4}{\Delta t^{2}}\sin^{2}(\frac{\pi}{2}\frac{j}{n+1}),$ $j=1,2,\cdots n$. Therefore the eigenvalues of $A+A^{T}$ are positive, hence the matrix is positive definite. Therefore the real part of the eigenvalues of $M=A+iB$ are strictly positive.
Remark 4.12.
[…]
Corollary 4.13.
Let $D^{\alpha}_{apx}u^{m}$ be the L1 approximation of Caputo derivative of $D^{\alpha}_{t}u(t_{m})$ on the temporal graded mesh $0=t_{0}<t_{1}<\cdots<t_{n}=T$, [8]. Then
[…]
$=\frac{1}{\Gamma(2-\alpha)}\sum_{j=0}^{m-1}\frac{u^{j+1}-u^{j}}{t_{j+1}-t_{j}}\left[(t_{m}-t_{j})^{1-\alpha}-(t_{m}-t_{j+1})^{1-\alpha}\right]$
$=\frac{1}{\Gamma(2-\alpha)}\left[d_{m,1}u^{m}-d_{m,m}u^{0}+\sum_{j=1}^{m-1}\left(d_{m,j+1}-d_{m,j}\right)u^{m-j}\right]$
choosing the initial condition $u^{0}=0$ and define
[…]
Now $(t_{m}-s)^{-\alpha}$ is monotonic increasing function for $s<t_{m}$ and $10<\alpha<$ 1. Hence by using integral mean value theorem one get $0<d_{m,j+1}<d_{m,j}$ for all $m$ and $0\leq j\leq m$. Therefore the difference matrix corresponding to the L1 scheme takes the form
$A=\begin{bmatrix}d_{1,1}&&&&\\ d_{2,2}-d_{2,1}&d_{2,1}&&&\\ \vdots&\ddots&\ddots&&\\ d_{n-1,n-1}-d_{n-1,n-2}&\cdots&&d_{n-1,1}&\\ d_{n,n}-d_{n,n-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}\end{bmatrix}.$
Clearly the $m^{th}$ row sum of absolute elements of matrix $A$ except the diagonal element is $R_{m}:=|(d_{m,2}-d_{m,1})|+\cdots+|(d_{m,m}-d_{m,m-1})|=d_{m,1}-d_{m,m}<d_{m,1}$. Hence by using Gershgorin circle theorem we can say that the eigenvalues of $A+A^{T}$ is strictly positive. Therefore using lemma 4.9 we can conclude that the real part of the eigenvalues of matrix $M=A+iB$, where $B$ takes the form as in corollary 4.11, are strictly positive.
Lemma 4.14.
Let $M=Q+P$ , where $M,Q,P$ are square matrices of order $n$ in $\mathbb{C}$. And $Q$ is diagonalizable, i.e. $Q=XDX^{-1}$. Then every eigenvalue of $M$ lies within the radius $R_{i}$, where $R_{i}$ is the Gershgorin radius of $R=X^{-1}PX$, centered at the addition of some eigenvalue of $Q$ and corresponding diagonal element of $R$.
Proof 4.15.
We have $M=Q+P$ and Q is diagonalizable. Therefore
$M$ $=XDX^{-1}+P$
$X^{-1}MX$ $=D+X^{-1}PX$
$=D+R.$
We know eigenvalues of $M$ and eigenvalues of $X^{-1}MX$ are same. Now we use Gershgorin circle theorem idea to prove the result. Let $\lambda$ be the eigenvalue of $X^{-1}MX$ associate with the eigenvector $y=(y_{j})$ and let $y_{i}$ is the largest absolute value of $y$. Then we have
$X^{-1}MXy$ $=\lambda y$
$(D+R)y$ $=\lambda y.$
Considering the $i^{th}$ equation from the system using $D=(d_{ij})_{n\times n}$ and $R=(r_{ij})_{n\times n}$ matrix elements we have
$\sum_{j=1}^{n}(d_{ij}+r_{ij})y_{j}$ $=\lambda y_{i}$
$d_{ii}y_{i}+\sum_{j=1}^{n}r_{ij}y_{j}$ $=\lambda y_{i}$
$\sum_{j=1,j\neq i}^{n}r_{ij}y_{j}$ $=(\lambda-d_{ii}-r_{ii})y_{i}$
Now using the triangular inequality we have
$|\lambda-d_{ii}-r_{ii}|\leq\sum_{j=1,j\neq i}^{n}|r_{ij}|\frac{|y_{j}|}{|y_{i}|}$
As $y_{i}$ is the largest absolute value in $y$ so
$|\lambda-d_{ii}-r_{ii}|$ $\leq\sum_{j=1,j\neq i}^{n}|r_{ij}|=R_{i}.$
The next result gives the estimate on Gershgorin radius $R_{i}$ for specific situation.
Lemma 4.16.
Let $M=S+J$, where $S$ and $J$ are respectively the lower triangular matrix and unit anti-diagonal matrix of order $n$. Then Greshgorin radius can be easily found out from the relation (14) for $n$ even and from the relation (15) for $n$ odd.
Proof 4.17.
[…]
(14) $X^{T}SX$ $=\frac{1}{2}\begin{bmatrix}I_{m,m}&J_{m,m}\\ J_{m,m}&-I_{m,m}\end{bmatrix}\begin{bmatrix}A_{m,m}&O_{m,m}\\ B_{m,m}&C_{m,m}\end{bmatrix}\begin{bmatrix}I_{m,m}&J_{m,m}\\ J_{m,m}&-I_{m,m}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}A_{m,m}+J_{m,m}B_{m,m}+J_{m,m}C_{m,m}J_{m,m}&A_{m,m}J_{m,m}-J_{m,m}C_{m,m}+J_{m,m}B_{m,m}J_{m,m}\\ J_{m,m}A_{m,m}-C_{m,m}J_{m,m}-B_{m,m}&C_{m,m}-B_{m,m}J_{m,m}+J_{m,m}A_{m,m}J_{m,m}\end{bmatrix}.$
[…]
(15) $X^{T}SX$
$=\frac{1}{2}\begin{bmatrix}I_{m,m}&O_{m,1}&J_{m,m}\\ O_{1,m}&\sqrt{2}I_{1,1}&O_{1,m}\\ J_{m,m}&O_{m,1}&-I_{m,m}\end{bmatrix}\begin{bmatrix}A_{m,m}&O_{m,1}&O_{m,m}\\ B_{1,m}&cI_{1,1}&O_{1,m}\\ C_{m,m}&D_{m,1}&E_{m,m}\end{bmatrix}\begin{bmatrix}I_{m,m}&O_{m,1}&J_{m,m}\\ O_{1,m}&\sqrt{2}I_{1,1}&O_{1,m}\\ J_{m,m}&O_{m,1}&-I_{m,m}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}A_{m,m}+J_{m,m}C_{m,m}+J_{m,m}E_{m,m}J_{m,m}&\sqrt{2}J_{m,m}D_{m,1}&A_{m,m}J_{m,m}-J_{m,m}E_{m,m}+J_{m,m}C_{m,m}J_{m,m}\\ \sqrt{2}B_{1,m}&2cI_{1,1}&\sqrt{2}B_{1,m}J_{m,m}\\ J_{m,m}A_{m,m}-E_{m,m}J_{m,m}-C_{m,m}&\sqrt{2}D_{m,1}&J_{m,m}A_{m,m}J_{m,m}-C_{m,m}J_{m,m}+E_{m,m}\end{bmatrix}.$
Corollary 4.18.
Let $S$ be the first order backward difference matrix so takes the same form as in 4.11. For $n$ even case: $A_{m,m}=C_{m,m}=I_{m,m}-L_{m,m}$, where $L_{m,m}$ is lower shift matrix. So,
$X^{T}SX$
$=\begin{bmatrix}I_{m,m}&O_{m,m}\\ O_{m,m}&I_{m,m}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}+J_{m,m}B_{m,m}&-L_{m,m}J_{m,m}+J_{m,m}L_{m,m}+J_{m,m}B_{m,m}J_{m,m}\\ -J_{m,m}L_{m,m}+L_{m,m}J_{m,m}-B_{m,m}&-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}-B_{m,m}J_{m,m}\end{bmatrix}.$
Now the matrix block takes the form:
$-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}+J_{m,m}B_{m,m}$ $=\begin{bmatrix}&-1&&&\\ -1&&-1&&\\ &\ddots&&\ddots&\\ &&-1&&-1\\ &&&-1&-1\end{bmatrix},$
$-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}-B_{m,m}J_{m,m}$ $=\begin{bmatrix}1&-1&&&\\ -1&&-1&&\\ &\ddots&&\ddots&\\ &&-1&&-1\\ &&&-1&\end{bmatrix},$
$-L_{m,m}J_{m,m}+J_{m,m}L_{m,m}+J_{m,m}B_{m,m}J_{m,m}$ $=\begin{bmatrix}&&&1&\\ &&1&&-1\\ &\iddots&&\iddots&\\ 1&&-1&&\\ -1&-1&&&\end{bmatrix},$
$-J_{m,m}L_{m,m}+L_{m,m}J_{m,m}-B_{m,m}$ $=\begin{bmatrix}&&&-1&1\\ &&-1&&1\\ &\iddots&&\iddots&\\ -1&&1&&\\ &1&&&\end{bmatrix}.$
[…]
$X^{T}SX$
$=\begin{bmatrix}I_{m,m}&O_{m,1}&O_{m,m}\\ O_{1,m}&I_{1.1}&O_{1,m}\\ O_{m,m}&O_{m,1}&I_{m,m}\end{bmatrix}+\frac{1}{2}\begin{bmatrix}-L_{m,m}-J_{m,m}L_{m,m}J_{m,m}&\sqrt{2}J_{m,m}D_{m,1}&-L_{m,m}J_{m,m}+J_{m,m}L_{m,m}\\ \sqrt{2}B_{1,m}&O_{1,1}&\sqrt{2}B_{1,m}J_{m,m}\\ -J_{m,m}L_{m,m}+L_{m,m}J_{m,m}&\sqrt{2}D_{m,1}&-J_{m,m}L_{m,m}J_{m,m}-L_{m,m}\end{bmatrix}.$
Now using the similar arguments as in the even case we can conclude that each Greshgorin radius $R_{i}\leq 2$.
Corollary 4.19.
Let $S$ be the Caputo finite difference matrix as in corollary 4.13 and $n$ be even. Then
$A_{m,m}$ $=\begin{bmatrix}d_{1,1}&&&&\\ d_{2,2}-d_{2,1}&d_{2,1}&&&\\ \vdots&&\ddots&&\\ d_{m-1,m-1}-d_{m-1,m-2}&\cdots&&d_{m-1,1}&\\ d_{m,m}-d_{m,m-1}&\cdots&&d_{m,2}-d_{m,1}&d_{m,1}\end{bmatrix},$
$C_{m,m}$ $=\begin{bmatrix}d_{m+1,1}&&&&\\ d_{m+2,2}-d_{m+2,1}&d_{m+2,1}&&&\\ \vdots&&\ddots&&\\ d_{n-1,m-1}-d_{n-1,m-2}&\cdots&&d_{n-1,1}&\\ d_{n,m}-d_{n,m-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}\end{bmatrix},$
$B_{m,m}$ $=\begin{bmatrix}d_{m+1,m+1}-d_{m+1,m}&&\cdots&&d_{m+1,2}-d_{m+1,1}\\ \vdots&&\vdots&&\vdots\\ d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+1}-d_{n,m}\end{bmatrix}.$
Therefore from the relation (14) different blocks of $X^{T}SX$ matrix are:
$A_{m,m}+J_{m,m}C_{m,m}J_{m,m}+J_{m,m}B_{m,m}$
$=\begin{bmatrix}d_{1,1}+d_{n,1}&d_{n,2}-d_{n,1}&&\cdots&d_{n,m}-d_{n,m-1}\\ d_{2,2}-d_{2,1}&d_{2,1}+d_{n-1,1}&&\cdots&d_{n-1,m-1}-d_{n-1,m-2}\\ \vdots&&\ddots&&\vdots\\ d_{m-1,m-1}-d_{m-1,m-2}&&\cdots&d_{m-1,1}+d_{m+2,1}&d_{m+2,2}-d_{m+2,1}\\ d_{m,m}-d_{m,m-1}&&\cdots&d_{m,2}-d_{m,1}&d_{m,1}+d_{m+1,1}\end{bmatrix}$
$+\begin{bmatrix}d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+1}-d_{n,m}\\ \vdots&&\vdots&&\vdots\\ d_{m+1,m+1}-d_{m+1,m}&&\cdots&&d_{m+1,2}-d_{m+1,1}\end{bmatrix},$
For computational simplicity we take the Gershgorin radius from this block for the center $\underline{C}_{1}=d_{1,1}+d_{n,1},\underline{C}_{2}=d_{2,1}+d_{n-1,1},\cdots,\underline{C}_{m}=d_{m,1}+d_{m+1,1}$ are
$\underline{R}_{1}$ $=|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|+|d_{n,m+1}-d_{n,m}|+\cdots+|d_{n,n}-d_{n,n-1}|$
$=d_{n,1}-d_{n,n},$
$\underline{R}_{2}$ $=|d_{2,2}-d_{2,1}|+\left(|d_{n-1,2}-d_{n-1,1}|+\cdots+|d_{n-1,n-1}-d_{n-1,n-2}|\right)$
$=(d_{2,1}-d_{2,2})+(d_{n-1,1}-d_{n-1,n-1}),$
$\vdots$
$\underline{R}_{m}$ $=(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|)+(|d_{m+1,2}-d_{m+1,1}|+\cdots+|d_{m+1,m+1}-d_{m+1,m}|)$
$=(d_{m,1}-d_{m,m})+(d_{m+1,1}-d_{m+1,m+1}).$
$C_{m,m}-B_{m,m}J_{m,m}+J_{m,m}A_{m,m}J_{m,m}$
$=\begin{bmatrix}d_{m+1,1}+d_{m,1}&d_{m,2}-d_{m,1}&&\cdots&d_{m,m}-d_{m,m-1}\\ d_{m+2,2}-d_{m+2,1}&d_{m+2,1}+d_{m-1,1}&&\cdots&d_{m-1,m-1}-d_{m-1,m-2}\\ \vdots&&\ddots&&\vdots\\ d_{n-1,m-1}-d_{n-1,m-2}&\cdots&&d_{n-1,1}+d_{2,1}&d_{2,2}-d_{2,1}\\ d_{n,m}-d_{n,m-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}+d_{1,1}\end{bmatrix}$
$-\begin{bmatrix}d_{m+1,2}-d_{m+1,1}&&\cdots&&d_{m+1,m+1}-d_{m+1,m}\\ \vdots&&\vdots&&\vdots\\ d_{n,m+1}-d_{n,m}&&\cdots&&d_{n,n}-d_{n,n-1}\end{bmatrix}$
Again for computational simplicity we take the Gershgorin radius from this block for the center $\underline{C}_{m+1}=d_{m+1,1}+d_{m,1},\underline{C}_{m+2}=d_{m+2,1}+d_{m-1,1},\cdots,\underline{C}_{n}=d_{n,1}+d_{1,1}$ are
$\underline{R}_{m+1}$ $=(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|)+(|d_{m+1,2}-d_{m+1,1}|+\cdots+|d_{m+1,m+1}-d_{m+1,m}|)$
$=(d_{m,1}-d_{m,m})+(d_{m+1,1}-d_{m+1,m+1}),$
$\vdots$
$\underline{R}_{n}$ $=|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|+|d_{n,m+1}-d_{n,m}|+\cdots+|d_{n,n}-d_{n,n-1}|$
$=d_{n,1}-d_{n,n}.$
$A_{m,m}J_{m,m}-J_{m,m}C_{m,m}+J_{m,m}B_{m,m}J_{m,m}$
$=\begin{bmatrix}-(d_{n,m}-d_{n,m-1})&\cdots&&-(d_{n,2}-d_{n,1})&d_{1,1}-d_{n,1}\\ -(d_{n-1,m-1}-d_{n-1,m-2})&\cdots&&d_{2,1}-d_{n-1,1}&d_{2,2}-d_{2,1}\\ &&\iddots&&\vdots\\ -(d_{m+2,2}-d_{m+2,1})&d_{m-1,1}-d_{m+2,1}&&\cdots&d_{m-1,m-1}-d_{m-1,m-2}\\ d_{m,1}-d_{m+1,1}&d_{m,2}-d_{m,1}&&\cdots&d_{m,m}-d_{m,m-1}\end{bmatrix}$
$+\begin{bmatrix}d_{n,m+1}-d_{n,m}&&\cdots&&d_{n,n}-d_{n,n-1}\\ \vdots&&\vdots&&\vdots\\ d_{m+1,2}-d_{m+1,1}&&\cdots&&d_{m+1,m+1}-d_{m+1,m}\end{bmatrix}$
As this is anti-diagonal block so the contribution for the Gershgorin radius from this block is:
$\bar{R}_{1}$ $=\left(|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|+|d_{n,m+1}-d_{n,m}|+\cdots+|d_{n,n}-d_{n,n-1}|\right)+|d_{1,1}-d_{n,1}|$
$=(d_{n,1}-d_{n,n})+|d_{1,1}-d_{n,1}|,$
$\bar{R}_{2}$ $=\left(|d_{n-1,2}-d_{n-1,1}|+\cdots+|d_{n-1,n-1}-d_{n-1,n-2}|\right)+|d_{2,1}-d_{n-1,1}|+|d_{2,2}-d_{2,1}|$
$=(d_{n-1,1}-d_{n-1,n-1})+|d_{2,1}-d_{n-1,1}|+(d_{2,1}-d_{2,2}),$
$\vdots$
$\bar{R}_{m}$ $=|d_{m,1}-d_{m+1,1}|+\left(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|\right)+\left(|d_{m+1,2}-d_{m+1,1}|+\cdots+|d_{m+1,m+1}-d_{m+1,m}|\right)$
$=(d_{m+1,1}-d_{m+1,m+1})+|d_{m,1}-d_{m+1,1}|+(d_{m,1}-d_{m,m}).$
$J_{m,m}A_{m,m}-C_{m,m}J_{m,m}-B_{m,m}$
$=\begin{bmatrix}d_{m,m}-d_{m,m-1}&\cdots&&d_{m,2}-d_{m,1}&d_{m,1}-d_{m+1,1}\\ d_{m-1,m-1}-d_{m-1,m-2}&\cdots&&d_{m-1,1}-d_{m+2,1}&-(d_{m+2,2}-d_{m+2,1})\\ &&\iddots&&\vdots\\ d_{2,2}-d_{2,1}&d_{2,1}-d_{n-1,1}&&\cdots&-(d_{n-1,m-1}-d_{n-1,m-2})\\ d_{1,1}-d_{n,1}&-(d_{n,2}-d_{n,1})&&\cdots&-(d_{n,m}-d_{n,m-1})\end{bmatrix}$
$-\begin{bmatrix}d_{m+1,m+1}-d_{m+1,m}&&\cdots&&d_{m+1,2}-d_{m+1,1}\\ \vdots&&\vdots&&\vdots\\ d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+1}-d_{n,m}\end{bmatrix}$
Again this is anti-diagonal block so the contribution for the Gershgorin radius from this block is:
$\bar{R}_{m+1}$ $=(d_{m,1}-d_{m,m})+|d_{m,1}-d_{m+1,1}|+(d_{m+1,1}-d_{m+1,m+1}),$
$\vdots$
$\bar{R}_{n}$ $=|d_{1,1}-d_{n,1}|+(d_{n,1}-d_{n,n}).$
Hence when $n$ is even the Gershgorin radius of the matrix $X^{T}SX$ for center $C_{1}=\frac{d_{1,1}+d_{n,1}}{2},\cdots,C_{m}=\frac{d_{m,1}+d_{m+1,1}}{2},C_{m+1}=\frac{d_{m+1,1}+d_{m,1}}{2},\cdots,C_{n}=\frac{d_{n,1}+d_{1,1}}{2}$ are
$R_{1}$ $=(d_{n,1}-d_{n,n})+\frac{1}{2}|d_{1,1}-d_{n,1}|\leq\frac{3}{2}(d_{1,1}+d_{n,1}),$
$R_{2}$ $=(d_{n-1,1}-d_{n-1,n-1})+\frac{1}{2}|d_{2,1}-d_{n-1,1}|+(d_{2,1}-d_{2,2})\leq\frac{3}{2}(d_{2,1}+d_{n-1,1}),$
$\vdots$
$R_{m}$ $=(d_{m+1,1}-d_{m+1,m+1})+\frac{1}{2}|d_{m,1}-d_{m+1,1}|+(d_{m,1}-d_{m,m})\leq\frac{3}{2}(d_{m,1}+d_{m+1,1}),$
$R_{m+1}$ $=(d_{m,1}-d_{m,m})+\frac{1}{2}|d_{m,1}-d_{m+1,1}|+(d_{m+1,1}-d_{m+1,m+1})\leq\frac{3}{2}(d_{m,1}+d_{m+1,1}),$
$\vdots$
$R_{n}$ $=\frac{1}{2}|d_{1,1}-d_{n,1}|+(d_{n,1}-d_{n,n})\leq\frac{3}{2}(d_{1,1}+d_{n,1}).$
When $n$ is odd and $m:=\frac{n-1}{2}$ from relation 15 we have:
$X^{T}SX$
$=\frac{1}{2}\begin{bmatrix}A_{m,m}+J_{m,m}C_{m,m}+J_{m,m}E_{m,m}J_{m,m}&\sqrt{2}J_{m,m}D_{m,1}&A_{m,m}J_{m,m}-J_{m,m}E_{m,m}+J_{m,m}C_{m,m}J_{m,m}\\ \sqrt{2}B_{1,m}&2cI_{1,1}&\sqrt{2}B_{1,m}J_{m,m}\\ J_{m,m}A_{m,m}-E_{m,m}J_{m,m}-C_{m,m}&\sqrt{2}D_{m,1}&J_{m,m}A_{m,m}J_{m,m}-C_{m,m}J_{m,m}+E_{m,m}\end{bmatrix}.$
and
$A_{m,m}$ $=\begin{bmatrix}d_{1,1}&&&&\\ d_{2,2}-d_{2,1}&d_{2,1}&&&\\ \vdots&&\ddots&&\\ d_{m-1,m-1}-d_{m-1,m-2}&\cdots&&d_{m-1,1}&\\ d_{m,m}-d_{m,m-1}&\cdots&&d_{m,2}-d_{m,1}&d_{m,1}\end{bmatrix},$
$E_{m,m}$ $=\begin{bmatrix}d_{m+2,1}&&&&\\ d_{m+3,2}-d_{m+3,1}&d_{m+3,1}&&&\\ \vdots&&\ddots&&\\ d_{n-1,m-1}-d_{n-1,m-2}&\cdots&&d_{n-1,1}&\\ d_{n,m}-d_{n,m-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}\end{bmatrix},$
$C_{m,m}$ $=\begin{bmatrix}d_{m+2,m+2}-d_{m+2,m+1}&&\cdots&&d_{m+2,3}-d_{m+2,2}\\ \vdots&&\vdots&&\vdots\\ d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+2}-d_{n,m+1}\end{bmatrix}.$
and $B_{1,m}=[(d_{m+1,m+1}-d_{m+1,m})\,\cdots\,(d_{m+1,2}-d_{m+1,1})]$, $c=d_{m+1,1}$, $D_{m,1}=[(d_{m+2,2}-d_{m+2,1})\,\cdots\,(d_{n,m+1}-d_{n,m})]^{T}$. Hence block of $X^{T}SX$ are:
$A_{m,m}+J_{m,m}C_{m,m}+J_{m,m}E_{m,m}J_{m,m}$
$=\begin{bmatrix}d_{1,1}+d_{n,1}&d_{n,2}-d_{n,1}&&\cdots&d_{n,m}-d_{n,m-1}\\ d_{2,2}-d_{2,1}&d_{2,1}+d_{n-1,1}&&\cdots&d_{n-1,m-1}-d_{n-1,m-2}\\ \vdots&&\ddots&&\vdots\\ d_{m-1,m-1}-d_{m-1,m-2}&&\cdots&d_{m-1,1}+d_{m+3,1}&d_{m+3,2}-d_{m+3,1}\\ d_{m,m}-d_{m,m-1}&&\cdots&d_{m,2}-d_{m,1}&d_{m,1}+d_{m+2,1}\end{bmatrix}$
$+\begin{bmatrix}d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+2}-d_{n,m+1}\\ \vdots&&\vdots&&\vdots\\ d_{m+2,m+2}-d_{m+2,m+1}&&\cdots&&d_{m+2,3}-d_{m+2,2}\end{bmatrix},$
For computational simplicity we take the Gershgorin radius from this block for the center $\underline{C}_{1}=d_{1,1}+d_{n,1},\underline{C}_{2}=d_{2,1}+d_{n-1,1},\cdots,\underline{C}_{m}=d_{m,1}+d_{m+2,1}$ are
$\underline{R}_{1}$ $=(|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|)+(|d_{n,m+2}-d_{n,m+1}|+\cdots+|d_{n,n}-d_{n,n-1}|)$
$=(d_{n,1}-d_{n,m})+(d_{n,m+1}-d_{n,n}),$
$\vdots$
$\underline{R}_{m}$ $=(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|)+(|d_{m+2,3}-d_{m+2,2}|+\cdots+|d_{m+2,m+2}-d_{m+2,m+1}|)$
$=(d_{m,1}-d_{m,m})+(d_{m+2,2}-d_{m+2,m+2}).$
$J_{m,m}A_{m,m}J_{m,m}-C_{m,m}J_{m,m}+E_{m,m}$
$=\begin{bmatrix}d_{m+2,1}+d_{m,1}&d_{m,2}-d_{m,1}&&\cdots&d_{m,m}-d_{m,m-1}\\ d_{m+3,2}-d_{m+3,1}&d_{m+3,1}+d_{m-1,1}&&\cdots&d_{m-1,m-1}-d_{m-1,m-2}\\ \vdots&&\ddots&&\vdots\\ d_{n-1,m-1}-d_{n-1,m-2}&\cdots&&d_{n-1,1}+d_{2,1}&d_{2,2}-d_{2,1}\\ d_{n,m}-d_{n,m-1}&\cdots&&d_{n,2}-d_{n,1}&d_{n,1}+d_{1,1}\end{bmatrix}$
$-\begin{bmatrix}d_{m+2,3}-d_{m+2,2}&&\cdots&&d_{m+2,m+2}-d_{m+2,m+1}\\ \vdots&&\vdots&&\vdots\\ d_{n,m+2}-d_{n,m+1}&&\cdots&&d_{n,n}-d_{n,n-1}\end{bmatrix}$
Again for computational simplicity we take the Gershgorin radius from this block for the center $\underline{C}_{m+2}=d_{m+2,1}+d_{m,1},\underline{C}_{m+3}=d_{m+3,1}+d_{m-1,1},\cdots,\underline{C}_{n}=d_{n,1}+d_{1,1}$ are
$\underline{R}_{m+2}$ $=(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|)+(|d_{m+2,3}-d_{m+2,2}|+\cdots+|d_{m+2,m+2}-d_{m+2,m+1}|)$
$=(d_{m,1}-d_{m,m})+(d_{m+2,2}-d_{m+2,m+2}),$
$\vdots$
$\underline{R}_{n}$ $=(|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|)+(|d_{n,m+2}-d_{n,m+1}|+\cdots+|d_{n,n}-d_{n,n-1}|)$
$=(d_{n,1}-d_{n,m})+(d_{n,m+1}-d_{n,n}).$
$A_{m,m}J_{m,m}-J_{m,m}E_{m,m}+J_{m,m}C_{m,m}J_{m,m}$
$=\begin{bmatrix}-(d_{n,m}-d_{n,m-1})&\cdots&&-(d_{n,2}-d_{n,1})&d_{1,1}-d_{n,1}\\ -(d_{n-1,m-1}-d_{n-1,m-2})&\cdots&&d_{2,1}-d_{n-1,1}&d_{2,2}-d_{2,1}\\ &&\iddots&&\vdots\\ -(d_{m+3,2}-d_{m+3,1})&d_{m-1,1}-d_{m+3,1}&&\cdots&d_{m-1,m-1}-d_{m-1,m-2}\\ d_{m,1}-d_{m+2,1}&d_{m,2}-d_{m,1}&&\cdots&d_{m,m}-d_{m,m-1}\end{bmatrix}$
$+\begin{bmatrix}d_{n,m+2}-d_{n,m+1}&&\cdots&&d_{n,n}-d_{n,n-1}\\ \vdots&&\vdots&&\vdots\\ d_{m+2,3}-d_{m+2,2}&&\cdots&&d_{m+2,m+2}-d_{m+2,m+1}\end{bmatrix}$
, As this is anti-diagonal block so the contribution for the Gershgorin radius from this block is:
$\bar{R}_{1}$ $=\left(|d_{n,2}-d_{n,1}|+\cdots+|d_{n,m}-d_{n,m-1}|\right)+\left(|d_{n,m+2}-d_{n,m+1}|+\cdots+|d_{n,n}-d_{n,n-1}|\right)+|d_{1,1}-d_{n,1}|$
$=(d_{n,1}-d_{n,m})+(d_{n,m+1}-d_{n,n})+|d_{1,1}-d_{n,1}|,$
$\vdots$
$\bar{R}_{m}$ $=|d_{m,1}-d_{m+2,1}|+\left(|d_{m,2}-d_{m,1}|+\cdots+|d_{m,m}-d_{m,m-1}|\right)+\left(|d_{m+2,3}-d_{m+2,2}|+\cdots+|d_{m+2,m+2}-d_{m+2,m+1}|\right)$
$=(d_{m+2,2}-d_{m+2,m+2})+|d_{m,1}-d_{m+2,1}|+(d_{m,1}-d_{m,m}).$
$J_{m,m}A_{m,m}-E_{m,m}J_{m,m}-C_{m,m}$
$=\begin{bmatrix}d_{m,m}-d_{m,m-1}&\cdots&&d_{m,2}-d_{m,1}&d_{m,1}-d_{m+2,1}\\ d_{m-1,m-1}-d_{m-1,m-2}&\cdots&&d_{m-1,1}-d_{m+3,1}&-(d_{m+3,2}-d_{m+3,1})\\ &&\iddots&&\vdots\\ d_{2,2}-d_{2,1}&d_{2,1}-d_{n-1,1}&&\cdots&-(d_{n-1,m-1}-d_{n-1,m-2})\\ d_{1,1}-d_{n,1}&-(d_{n,2}-d_{n,1})&&\cdots&-(d_{n,m}-d_{n,m-1})\end{bmatrix}$
$-\begin{bmatrix}d_{m+2,m+2}-d_{m+2,m+1}&&\cdots&&d_{m+2,3}-d_{m+2,2}\\ \vdots&&\vdots&&\vdots\\ d_{n,n}-d_{n,n-1}&&\cdots&&d_{n,m+2}-d_{n,m+1}\end{bmatrix},$
Again this is anti-diagonal block so the contribution for the Gershgorin radius from this block are:
$\bar{R}_{m+2}$ $=(d_{m,1}-d_{m,m})+|d_{m,1}-d_{m+2,1}|+(d_{m+2,2}-d_{m+2,m+2}),$
$\vdots$
$\bar{R}_{n}$ $=|d_{1,1}-d_{n,1}|+(d_{n,1}-d_{n,m})+(d_{n,m+1}-d_{n,n}).$
$J_{m,m}D_{m,1}=[(d_{n,m+1}-d_{n,m})\,\cdots\,(d_{m+2,2}-d_{m+2,1})]^{T}$, $B_{1,m}J_{m,m}=[(d_{m+1,2}-d_{m+1,1})\,\cdots\,(d_{m+1,m+1}-d_{m+1,m})]$. Hence when $n$ is odd the Gershgorin radius of the matrix $X^{T}SX$ for center $C_{1}=\frac{d_{1,1}+d_{n,1}}{2},\cdots,C_{m}=\frac{d_{m,1}+d_{m+2,1}}{2},C_{m+1}=d_{m+1,1},C_{m+2}=\frac{d_{m+2,1}+d_{m,1}}{2},\cdots,C_{n}=\frac{d_{n,1}+d_{1,1}}{2}$ are
$R_{1}$ $=(d_{n,1}-d_{n,m})+(d_{n,m+1}-d_{n,n})+\frac{1}{\sqrt{2}}(d_{n,m+1}-d_{n,m})+\frac{1}{2}|d_{1,1}-d_{n,1}|\leq\frac{3}{2}(d_{1,1}+d_{n,1}),$
$\vdots$
$R_{m}$ $=(d_{m+2,2}-d_{m+2,m+2})+\frac{1}{\sqrt{2}}(d_{m+2,2}-d_{m+2,1})+\frac{1}{2}|d_{m,1}-d_{m+2,1}|+(d_{m,1}-d_{m,m})\leq\frac{3}{2}(d_{m,1}+d_{m+2,1}),$
$R_{m+1}$ $=\frac{1}{\sqrt{2}}(|d_{m+1,m+1}-d_{m+1,m}|+\cdots+|d_{m+1,2}-d_{m+1,1}|)=\frac{1}{\sqrt{2}}(d_{m+1,1}-d_{m+1,m+1})\leq\frac{1}{\sqrt{2}}d_{m+1,1},$
$R_{m+2}$ $=(d_{m,1}-d_{m,m})+\frac{1}{2}|d_{m,1}-d_{m+1,1}|+(d_{m+2,2}-d_{m+2,m+2})+\frac{1}{\sqrt{2}}(d_{m+2,2}-d_{m+2,1})\leq\frac{3}{2}(d_{m,1}+d_{m+2,1}),$
$\vdots$
$R_{n}$ $=\frac{1}{2}|d_{1,1}-d_{n,1}|+(d_{n,1}-d_{n,m})+(d_{n,m+1}-d_{n,n})+\frac{1}{\sqrt{2}}(d_{n,m+1}-d_{n,m})\leq\frac{3}{2}(d_{1,1}+d_{n,1}).$
Lemma 4.20.
Let $L=\begin{bmatrix}L_{1}&O\\ L_{2}&l\end{bmatrix}$ be a real lower triangular matrix of order $n+1$ and $L_{1}$ block is of order $n$. $B=\begin{bmatrix}B_{1}&O\\ O&0\end{bmatrix}$ be a real block matrix of order $n+1$ with block $B_{1}$ of order $n$ and $O$ be the zero matrix. If $(L_{1}-lI+\frac{i}{\sqrt{\sigma}}B_{1})$ is non-singular then the matrix $L+\frac{i}{\sqrt{\sigma}}B$ is similar to matrix $\bar{L}=\begin{bmatrix}L_{1}+\frac{i}{\sqrt{\sigma}}B_{1}&O\\ O&l\end{bmatrix}$.
Proof 4.21.
Let $X=\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}$ be the block matrix of order $n+1$, where $X_{11}$ is of order $n$ block. Let
$\begin{bmatrix}L_{1}+\frac{i}{\sqrt{\sigma}}D_{A}&0\\ L_{2}&l\end{bmatrix}\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}$ $=\begin{bmatrix}X_{11}&X_{12}\\ X_{21}&X_{22}\end{bmatrix}\begin{bmatrix}L_{1}+\frac{i}{\sqrt{\sigma}}D_{A}&0\\ 0&l\end{bmatrix}$
$\begin{bmatrix}(L_{1}+\frac{i}{\sqrt{\sigma}}D_{A})X_{11}&(L_{1}+\frac{i}{\sqrt{\sigma}}D_{A})X_{12}\\ L_{2}X_{11}+lX_{21}&L_{2}X_{12}+lX_{22}\end{bmatrix}$ $=\begin{bmatrix}X_{11}(L_{1}+\frac{i}{\sqrt{\sigma}}D_{A})&X_{12}l\\ X_{21}(L_{1}+\frac{i}{\sqrt{\sigma}}D_{A})&X_{22}l\end{bmatrix}$
We choose $X_{11}=I,X_{12}=0,X_{22}=I$ then we have $X_{21}(L_{1}+\frac{i}{\sqrt{\sigma}}B_{1})=lX_{21}+L_{2}$. Therefore the system has solution if the matrix $(L_{1}-lI+\frac{i}{\sqrt{\sigma}}B_{1})$ is invertiable. Let $\underline{L}=(L_{1}-lI+\frac{i}{\sqrt{\sigma}}B_{1})^{-1}$ exist from our assumption. Therefor the matrix $L+\frac{i}{\sqrt{\sigma}}B$ is similar to $\bar{L}$ by the matrix $\begin{bmatrix}I&0\\ L_{2}+lL_{2}\underline{L}&1\end{bmatrix}$.
Remark 4.22.
Most of the practical cases $\sigma$ is very small therefore we can use lemma 4.16 and its corollaries 4.18 and 4.19 to show that $(L_{1}-lI+\frac{i}{\sqrt{\sigma}}B_{1})$ is invertiable when $B_{1}$ be the unit anti-diagonal matrix $J$. Again if $L_{1}-lI$ is toeplitz and $B_{1}=J$ then $|\det(L_{1}-lI+\frac{i}{\sqrt{\sigma}}J)|=|\det((L_{1}-lI)J+\frac{i}{\sqrt{\sigma}}I)|$. Now $(L_{1}-lI)J$ is symmetric so eigenvalues are all real hence $\det((L_{1}-lI)J+\frac{i}{\sqrt{\sigma}}I)\neq 0$ for any $\sigma$.
Before going to main convergence results for DNWR and NNWR algorithm for the control problem we want to introduce a different kind of mesh grading technique.
5 Mesh Grading
We know that for fractional diffusion graded mesh in time give higher order accuracy than the uniform mesh, therefore one can consider as usual one sided graded mesh [9]. Now in control problem we need to solve both state and adjoint problem so the obvious extension of one sided graded mesh to both sided grading. Let $[0,T]$ be the total time domain and it is discretize into $2n:n\in\mathcal{N}$ non-overlapping sub-intervals by the nodes $t_{0}<t_{1}<\cdots<t_{2n}$. Now first we take usual graded mesh for time window $[0,T/2]$ using the nodes $t_{0}<t_{1}<\cdots<t_{n}$. Now for the remaining time window $[T/2,T]$ we take the similar grading but started from right end i.e. from $T$ using the nodes $t_{n+1}<t_{n+2}<\cdots<t_{2n}$. Hence for $j=0,1,\cdots,n$ time nodes $t_{j}=(j/n)^{(2-\alpha)/\alpha}T/2$ and $t_{2n-j}=T-t_{j}$.
6 Convergence of DNWR Algorithm
Now we analyze the convergence estimates of the DNWR algorithm (5)-(6), opting for the 1D model of the sub-diffusion and diffusion-wave equation to achieve better theoretical clarity and algebraic simplicity in our estimates. Our analysis focuses on a heterogeneous problem with $\kappa(\boldsymbol{x},t)=\kappa_{1}$ on $\Omega_{1}=(-h_{1},0)$ and $\kappa(\boldsymbol{x},t)=\kappa_{2}$ on $\Omega_{2}=(0,h_{2})$. We introduce the interface error $\omega^{(k)}(t)$ on state equation and $\upsilon^{(k)}(t)$ on adjoint equation and using the lemma 4.1 for transforming the Riemann-Liouville derivative to Caputo derivative the error euqation becomes:
(16) $\begin{array}[]{rcll}\begin{cases}{}_{0}D_{t}^{\alpha}y_{1}^{(k)}=\kappa_{1}\partial_{xx}y_{1}^{(k)}-\frac{p_{1}^{(k)}}{\sigma},&\textrm{in}\;(-h_{1},0),\\ y_{1}^{(k)}(\boldsymbol{x},0)=0,&\textrm{in}\;(-h_{1},0),\\ y_{1}^{(k)}(0,t)=\omega^{(k-1)}(t),\\ y_{1}^{(k)}(-h_{1},t)=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}{}_{0}D_{t}^{\alpha}y_{2}^{(k)}=\kappa_{1}\partial_{xx}y_{2}^{(k)}-\frac{p_{2}^{(k)}}{\sigma},&\textrm{in}\;(0,h_{2}),\\ y_{2}^{(k)}(\boldsymbol{x},0)=0,&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}y_{2}^{(k)}(0,t)=\kappa_{1}\partial_{x}y_{1}^{(k)}(0,t),\\ y_{2}^{(k)}(h_{2},t)=0,\end{cases}\end{array}$
(17) $\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{1}^{(k)}=\kappa_{1}\partial_{xx}p_{1}^{(k)}+y_{1}^{(k)},&\textrm{in}\;(-h_{1},0),\\ p_{1}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;(-h_{1},0),\\ p_{1}^{(k)}(0,t)=\upsilon^{(k-1)}(t),\\ p_{1}^{(k)}(-h_{1},t)=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{2}^{(k)}=\kappa_{2}\partial_{xx}p_{2}^{(k)}+y_{2}^{(k)},&\textrm{in}\;(0,h_{2}),\\ p_{2}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}p_{2}^{(k)}(0,t)=\kappa_{1}\partial_{x}p_{1}^{(k)}(0,t),\\ p_{2}^{(k)}(h_{2},0)=0,\end{cases}\end{array}$
with the update step:
(18) $\omega^{(k)}(t)$ $=\theta y_{2}^{(k)}(0,t)+(1-\theta)\omega^{(k-1)}(t),$
(19) $\upsilon^{(k)}(t)$ $=\phi p_{2}^{(k)}(0,t)+(1-\phi)\upsilon^{(k-1)}(t),$
Now we will analyze the system using semi-discrete analysis. We take the time nodes $0=t_{0}<t_{1}<\cdots<t_{n+1}=T$, $n\in\mathbb{N}$. Now we may take graded or uniform time mesh. Now if we use L1 scheme for left and right sided Caputo fractional derivative then for uniform and both sided graded mesh time discretize left and right fractional derivative matrix $A_{t}$ and $\bar{A}_{t}$ will be same for any $n$ in the case of uniform mesh and odd $n$ for both sided graded mesh. But for usual one sided graded mesh they will be different. This is because for uniform and both sided graded mesh follow the reflection symmetry i.e.$t_{j}=T-t_{n+1-j}$ for $j=0,1,\cdots,n+1$. This condition is violated for one sided graded mesh. As graded mesh itself take care of the fractional order so as the the fractional order $\alpha$ tends to $1$, the graded mesh also tends to the uniform mesh. With these mechanism of discretization we have
$A_{t}:=\begin{bmatrix}d_{1,1}&&&&\\ d_{2,2}-d_{2,1}&d_{2,1}&&&\\ \vdots&\ddots&\ddots&&\\ d_{n,n}-d_{n,n-1}&\cdots&&d_{n,1}&\\ d_{n+1,n+1}-d_{n+1,n}&\cdots&&d_{n+1,2}-d_{n+1,1}&d_{n+1,1}\end{bmatrix}.\textit{ And }B_{t}:=\begin{bmatrix}&&&1&\\ &&1&&\\ &\iddots&&\\ 1&&&&\\ &&&&\end{bmatrix}$
be the order $n+1$ square matrix coming from the source term in the right hand side. Here $d_{m,j}$ for $11\leq j\leq m\leq n+$ 1 are given in (13). For the sake of the analysis we choose only the uniform and both sided graded mesh and have the following state and adjoint equation:
(20) $\begin{array}[]{rcll}\begin{cases}A_{t}Y_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}Y_{1}^{(k)}(\boldsymbol{x})-\frac{1}{\sigma}B_{t}P_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ Y_{1}^{(k)}(0)=\Omega^{(k-1)},\\ Y_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}A_{t}Y_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}Y_{2}^{(k)}(\boldsymbol{x})-\frac{1}{\sigma}B_{t}P_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}Y_{2}^{(k)}(0)=\kappa_{1}\partial_{x}Y_{1}^{(k)}(0),\\ Y_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
and
(21) $\begin{array}[]{rcll}\begin{cases}A_{t}P_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}P_{1}^{(k)}(\boldsymbol{x})+B_{t}Y_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ P_{1}^{(k)}(0)=\Upsilon^{(k-1)},\\ P_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}A_{t}P_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}P_{2}^{(k)}(\boldsymbol{x})+B_{t}Y_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}P_{2}^{(k)}(0)=\kappa_{1}\partial_{x}P_{1}^{(k)}(0),\\ P_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
with the update in discretize form:
(22) $\Omega^{(k)}$ $=\theta Y_{2}^{(k)}+(1-\theta)\Omega^{(k-1)},$
(23) $\Upsilon^{(k)}$ $=\phi P_{2}^{(k)}+(1-\phi)\Upsilon^{(k-1)},$
Here we have taken the discretization of $Y_{i}=[y_{i}^{1},y_{i}^{2},...,y_{i}^{n+1}]^{T}$ and $P_{i}$ into reverse direction i.e. $P_{i}=[p_{i}^{n},p_{i}^{n-1},...,p_{i}^{0}]^{T}$. For sake of analysis we combine both state and adjoint system into a single system and choosing the dependent variable $X_{i}=[Y_{i},P_{i}]^{T}$ Using the above two equations we form the system:
(24) $\begin{array}[]{rcll}\begin{cases}AX_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}X_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ X_{1}^{(k)}(0)=\Pi^{(k-1)},\\ X_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}AX_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}X_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}X_{2}^{(k)}(0)=\kappa_{1}\partial_{x}X_{1}^{(k)}(0),\\ X_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
Where, $A=\begin{bmatrix}A_{t}&\frac{1}{\sigma}B_{t}\\ -B_{t}&A_{t}\end{bmatrix}$. We choose $\Pi^{(k)}=[\Omega^{(k)},\Upsilon^{(k)}]^{T}$ for the combine form of the update condition for the new system and get:
(25) $\Pi^{(k)}=\begin{bmatrix}\theta I&0\\ 0&\phi I\end{bmatrix}X_{2}^{(k)}+\begin{bmatrix}(1-\theta)I&0\\ 0&(1-\phi)I\end{bmatrix}\Pi^{(k-1)}.$
Before going to the main convergent result of DNWR algorithm we first need to check that the time discretization is a convergent scheme of the continuous version, i.e. we need to check the consistency and stability of the numerical scheme. Which is achieved by the following theorem.
Theorem 6.1.
For the L1 numerical scheme in time fractional derivative the semi-discrete ode system in (24) is convergent for the pde system (16) and (17).
Proof 6.2.
We need to show that the numerical scheme is consistent and stable for the convergence of the solution. In discretization we have used L1-scheme on graded mesh or on uniform mesh, which is consistent. Now to show this scheme is stable we use the similar technique as in lemma 4.3. For that we first need the error equation of the semi discrete pde, which can be obtained by the substruction of the semi discrete pde with the continuous pde with the Taylor approximation of the time derivative. which is given as follows:
(26) $\begin{cases}AE_{1}(x)=\partial_{xx}E_{1}(x)+F(x),&\textrm{in}\;(-h_{1},0),\\ E_{1}(0)=0,\\ E_{1}(-h_{1})=0,\end{cases}$
Where $F(x)$ is the remaining term of the Taylor series expansion in time. To show the stability we use the Fourier frequency in space. We see that Fourier sine expansion is enough for that purpose. We consider each frequency separately and we have:
(27) $A\hat{E}_{1}=-\xi^{2}\hat{E}_{1}+\hat{F},$
$\hat{E}_{1}=(A+\xi^{2}I)^{-1}\hat{F}.$
Therefor the error depends on the absolute minimum eigenvalue of the matrix $(A+\xi^{2}I)$, which implies the eigenvalue of matrix $A$. Now $A=\begin{bmatrix}A_{t}&\frac{1}{\sigma}B_{t}\\ -B_{t}&A_{t}\end{bmatrix}$. From lemma 4.5 we know that eigenvalues of $A$ are the collection of eigenvalues of $A_{t}+\frac{i}{\sqrt{\sigma}}B_{t}$ and $A_{t}-\frac{i}{\sqrt{\sigma}}B_{t}$. Let, $A_{t}=\begin{bmatrix}A1_{t}&0\\ A2_{t}&a_{t}\end{bmatrix}$ be the order $n$ block matrix with $A1_{t}$ be the lower triangular block of order $(n-1)$ and $B_{t}=\begin{bmatrix}B1_{t}&0\\ 0&0\end{bmatrix}$ be the order $n$ block matrix with $B1_{t}$ be the anti-diagonal matrix of order $(n-1)$. From corollary 4.13 we know that each diagonal element from L1 scheme is positive, so $a_{t}\neq 0$. Again from the same corollary we have the matrix $A1_{t}$ is positive definite, hence from lemma 4.9 we conclude that the eigenvalues of $A_{t}+\frac{i}{\sqrt{\sigma}}B_{t}$ consist positive real part. In a similar manner we can also show the same for $A_{t}-\frac{i}{\sqrt{\sigma}}B_{t}$. Hence $(A+\xi^{2}I)$ will be a nonsingular matrix and the error will be bounded. Similar procedure can be used for the other part of (24) with Neumann boundary condition, where we have to use cosine function with certain phase difference.
Let, $A$ matrix is diagonalizable and $A=P\Lambda P^{-1}$. Solutions of (24) take the form:
$X_{1}^{(k)}(x)$ $=P\sinh\left((x+h_{1})\sqrt{\Lambda/\kappa_{1}}\right)\operatorname{cosech}(h_{1}\sqrt{\Lambda/\kappa_{1}})P^{-1}\Pi^{(k-1)},$
(28) $X_{2}^{(k)}(x)$ $=\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}P\sinh((x-h_{2})\sqrt{\Lambda/\kappa_{2}})\operatorname{sech}(h_{2}\sqrt{\Lambda/\kappa_{2}})\coth(h_{1}\sqrt{\Lambda/\kappa_{1}})P^{-1}\Pi^{(k-1)}.$
Substituting (28) into (25), the recurrence relation for $k\in\mathbb{N}$ becomes:
(29) $\Pi^{(k)}$ $=-\begin{bmatrix}\theta I&0\\ 0&\phi I\end{bmatrix}\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}P\tanh(h_{2}\sqrt{\Lambda/\kappa_{2}})\coth(h_{1}\sqrt{\Lambda/\kappa_{1}})P^{-1}\Pi^{(k-1)}+\begin{bmatrix}(1-\theta)I&0\\ 0&(1-\phi)I\end{bmatrix}\Pi^{(k-1)}.$
To ease the notional difficulties, define $a:=h_{1}/\sqrt{\kappa_{1}}\text{ and }b:=h_{2}/\sqrt{\kappa_{2}}$ for DNWR case. We first choose the most obvious case $a\neq b$, and define
(30) $\Pi^{(k)}$ $=-\begin{bmatrix}\theta I&0\\ 0&\phi I\end{bmatrix}\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}P\left(\tanh(b\sqrt{\Lambda})\coth(a\sqrt{\Lambda})-I\right)P^{-1}\Pi^{(k-1)}$
$+\begin{bmatrix}(1-(1+\sqrt{\kappa_{1}/\kappa_{2}})\theta)I&0\\ 0&(1-(1+\sqrt{\kappa_{1}/\kappa_{2}})\phi)I\end{bmatrix}\Pi^{(k-1)}.$
Choose $\theta=\phi=1/(1+\sqrt{\kappa_{1}/\kappa_{2}})$ we have:
(31) $\Pi^{(k)}$ $=-\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}P\left(\tanh(b\sqrt{\Lambda})\coth(a\sqrt{\Lambda})-I\right)P^{-1}\Pi^{(k-1)}.$
Theorem 6.3.
If for every eigenvalue $\lambda_{i}$ of $A$, the DNWR algorithm for optimal control converge if the quantity
$\left|\frac{\sinh((b-a)\sqrt{\lambda_{i}})}{\sinh(a\sqrt{\lambda_{i}})\cosh(b\sqrt{\lambda_{i}})}\right|<\frac{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}{\sqrt{\kappa_{1}}}.$
Lemma 6.4.
For any $\alpha,\beta$ nonzero positive real number and $z\in\mathbb{C}$
$\left|\frac{\exp(-\alpha z)}{1-\exp(-\beta z)}\right|\leq\frac{\exp(-\alpha\Re(z))}{1-\exp(-\beta\Re(z))}.$
Proof 6.5.
Let $z=x+iy$. Then
$\left|\frac{\exp(-\alpha z)}{1-\exp(-\beta z)}\right|$ $=\frac{\exp(-\alpha x)}{\sqrt{(1-\exp(-\beta x)\cos(\beta y))^{2}+(\exp(-\beta x)\sin(\beta y))^{2}}}$
$\leq\frac{\exp(-\alpha x)}{\sqrt{1-2\exp(-\beta x)\cos(\beta y)+\exp(-2\beta x)}}$
$\leq\frac{\exp(-\alpha x)}{1-\exp(-\beta x)}.$
Theorem 6.6.
Let $\lambda_{i}$ are the eigenvalues and $P$ is the eigenvector matrix of matrix $A$. If $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$, then the interface error term of the DNWR algorithm satisfies the following bound:
$\|\Pi^{(k)}\|_{\infty}<\left(\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\right)^{k}\left(\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}\right)^{k}\|P\|_{\infty}\|P^{-1}\|_{\infty}\|\Pi^{(0)}\|_{\infty}.$
Proof 6.7.
Left multiplying with $P^{-1}$ in (31) error equation becomes:
(32) $P^{-1}\Pi^{(k)}$ $=-\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\left(\tanh(b\sqrt{\Lambda})\coth(a\sqrt{\Lambda})-I\right)P^{-1}\Pi^{(k-1)}$
$=-\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\frac{\sinh((b-a)\sqrt{\Lambda})}{\sinh(b\sqrt{\Lambda})\cosh(a\sqrt{\Lambda})}P^{-1}\Pi^{(k-1)}$
Now consider the $i^{th}$ entry in (32) we have:
(33) $\left(P^{-1}\Pi^{(k)}\right)_{i}$ $=-\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\frac{\sinh((b-a)\sqrt{\lambda_{i}})}{\sinh(b\sqrt{\lambda_{i}})\cosh(a\sqrt{\lambda_{i}})}\left(P^{-1}\Pi^{(k-1)}\right)_{i}$
Let $(W)_{ii}:=\frac{\sinh((b-a)\sqrt{\lambda_{i}})}{\sinh(b\sqrt{\lambda_{i}})\cosh(a\sqrt{\lambda_{i}})}$ be the weight. And we will estimate the weight by simplifying the hyperbolic terms in terms of exponential and we have:
(34) $(W)_{ii}$ $=2\frac{\left(\exp((b-a)\sqrt{\lambda_{i}})-\exp(-(b-a)\sqrt{\lambda_{i}})\right)}{\left(\exp(a\sqrt{\lambda_{i}})-\exp(-a\sqrt{\lambda_{i}})\right)\left(\exp(b\sqrt{\lambda_{i}})+\exp(-b\sqrt{\lambda_{i}})\right)}$
$=2\frac{\left(\exp(-2a\sqrt{\lambda_{i}})-\exp(-2b\sqrt{\lambda_{i}})\right)}{\left(1-\exp(-2a\sqrt{\lambda_{i}})\right)\left(1+\exp(-2b\sqrt{\lambda_{i}})\right)}$
$=2\frac{\exp(-2a\sqrt{\lambda_{i}})-\exp(-2b\sqrt{\lambda_{i}})}{\left(1-\exp(-2a\sqrt{\lambda_{i}})\right)\left(1+\exp(-2b\sqrt{\lambda_{i}})\right)}$
$=\frac{2\exp(-2a\sqrt{\lambda_{i}})}{\left(1-\exp(-2a\sqrt{\lambda_{i}})\right)\left(1+\exp(-2b\sqrt{\lambda_{i}})\right)}-\frac{2\exp(-2b\sqrt{\lambda_{i}})}{\left(1-\exp(-2a\sqrt{\lambda_{i}})\right)\left(1+\exp(-2b\sqrt{\lambda_{i}})\right)}$
$:=I_{1}-I_{2}$
Now
(35) $|I_{1}|$ $=2\left|\frac{\exp(-a\sqrt{\lambda_{i}})}{\left(1-\exp(-2a\sqrt{\lambda_{i}})\right)}\right|\left|\frac{\exp(-a\sqrt{\lambda_{i}})}{\left(1+\exp(-2b\sqrt{\lambda_{i}})\right)}\right|$
Using the inequality from lemma 6.4, we have
$|I_{1}|$ $\leq 2\frac{\exp(-2a\Re(\sqrt{\lambda_{i}}))}{\left(1-\exp(-2a\Re(\sqrt{\lambda_{i}}))\right)\left(1-\exp(-2b\Re(\sqrt{\lambda_{i}}))\right)}$
Further using $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$, we obtain
$|I_{1}|$ $\leq 2\frac{\exp(-2a\lambda))}{\left(1-\exp(-2a\lambda))\right)\left(1-\exp(-2b\lambda))\right)}$
By similar argument
(36) $|I_{2}|$ $=2\left|\frac{\exp(-b\sqrt{\lambda_{i}})}{\left(1-\exp(-2a\sqrt{\lambda_{i}})\right)}\right|\left|\frac{\exp(-b\sqrt{\lambda_{i}})}{\left(1+\exp(-2b\sqrt{\lambda_{i}})\right)}\right|$
$\leq 2\frac{\exp(-2b\lambda)}{\left(1-\exp(-2a\lambda)\right)\left(1-\exp(-2b\lambda)\right)}$
Therefore
(37) $|(W)_{ii}|$ $\leq 2\frac{\exp(-2a\lambda)+\exp(-2b\lambda)}{\left(1-\exp(-2a\lambda)\right)\left(1-\exp(-2b\lambda)\right)}$
$\leq\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}$
Hence $\|W\|_{\infty}\leq\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}$. Now taking the supremum norm on both side in (37) we have:
(38) $\left\|P^{-1}\Pi^{(k)}\right\|_{\infty}$ $\leq\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}\left\|P^{-1}\Pi^{(k-1)}\right\|_{\infty}$
$\leq\left(\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\right)^{k}\left(\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}\right)^{k}\left\|P^{-1}\Pi^{(0)}\right\|_{\infty}$
Again using the norm inequality we have the required result:
$\|\Pi^{(k)}\|_{\infty}\leq\left(\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\right)^{k}\left(\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}\right)^{k}\|P\|_{\infty}\|P^{-1}\|_{\infty}\|\Pi^{(0)}\|_{\infty}.$
7 Convergence of NNWR Algorithm
To study the convergence of the NNWR algorithm (LABEL:NNWRD)-(11) in 1D, we first divide the spatial interval $(-L_{1},L_{2})$ into $N$ subintervals $(x_{i-1},x_{i})$ with no overlap or cross point, where $i=1,2,...,N$. We define $a_{i}$ as the length of each subinterval, and we assume different constant diffusion coefficients $\kappa_{i}$ in each subinterval $\Omega_{i}$. We also define $\omega_{i}^{(k)}(t)$ as the initial error along the artificial boundary at ${x_{i}}_{i=1}^{N-1}$ and set $\omega_{0}^{(k)}(t)=0$ at $x=x_{0}$ and $\omega_{N}^{(k)}(t)=0$ at $x=x_{N}$ for the original boundaries. By applying the Laplace transform in time, we can simplify the equations (LABEL:NNWRD)-(11) to
(39) $\begin{array}[]{rcll}\begin{cases}{}_{0}D_{t}^{\alpha}y_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}y_{i}^{(k)}-\frac{p_{i}^{(k)}}{\sigma},&\textrm{in}\;(x_{i-1},x_{i}),\\ y_{i}^{(k)}(\boldsymbol{x},0)=0,&\textrm{in}\;(x_{i-1},x_{i}),\\ y_{i}^{(k)}=\omega_{i-1}^{(k-1)},&\textrm{at }x_{i-1},i\neq 1,\\ y_{i}^{(k)}=\omega_{i}^{(k-1)},&\textrm{at }x_{i},i\neq N,\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}{}_{0}D_{t}^{\alpha}\psi_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}\psi_{i}^{(k)}-\frac{\chi_{i}^{(k)}}{\sigma},&\textrm{in}\;(x_{i-1},x_{i}),\\ \psi_{i}^{(k)}(\boldsymbol{x},0)=0,&\textrm{in}\;(x_{i-1},x_{i}),\\ -\kappa_{i}\frac{d}{dx}\psi_{i}^{(k)}=\kappa_{i-1}\frac{d}{dx}y_{i-1}^{(k)}-\kappa_{i}\frac{d}{dx}y_{i}^{(k)}&\textrm{at }x_{i-1},i\neq 1,\\ \kappa_{i}\frac{d}{dx}\psi_{i}^{(k)}=\kappa_{i}\frac{d}{dx}y_{i}^{(k)}-\kappa_{i+1}\frac{d}{dx}y_{i+1}^{(k)},&\textrm{at }x_{i},i\neq N-1.\end{cases}\end{array}$
(40) $\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}p_{i}^{(k)}+y_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ p_{i}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;(x_{i-1},x_{i}),\\ p_{i}^{(k)}=\upsilon_{i-1}^{(k-1)},&\textrm{at }x_{i-1},i\neq 1,\\ p_{i}^{(k)}=\upsilon_{i}^{(k-1)},&\textrm{at }x_{i},i\neq N,\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}\chi_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}\chi_{i}^{(k)}+\psi_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ \chi_{i}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;(x_{i-1},x_{i}),\\ -\kappa_{i}\frac{d}{dx}\chi_{i}^{(k)}=\kappa_{i-1}\frac{d}{dx}p_{i-1}^{(k)}+\kappa_{i}\frac{d}{dx}p_{i}^{(k)},&\textrm{at }x_{i-1},i\neq 1,\\ \kappa_{i}\frac{d}{dx}\chi_{i}^{(k)}=\kappa_{i}\frac{d}{dx}p_{i}^{(k)}+\kappa_{i+1}\frac{d}{dx}p_{i+1}^{(k)},&\textrm{at }x_{i},i\neq N.\end{cases}\end{array}$
The formula is then used to update the interface values for $i=1,\ldots,N-1$ becomes:
(41) $\omega_{i}^{(k)}(x_{i},t)$ $=\omega_{i}^{(k-1)}(x_{i},t)-\theta_{i}\left(\psi_{i}^{(k)}(x_{i},t)+\psi_{i+1}^{(k)}(x_{i},t)\right),$
(42) $\upsilon_{i}^{(k)}(x_{i},t)$ $=\upsilon_{i}^{(k-1)}(x_{i},t)-\phi_{i}\left(\chi_{i}^{(k)}(x_{i},t)+\chi_{i+1}^{(k)}(x_{i},t)\right).$
After discretization the system becomes:
(43) $\begin{array}[]{rcll}\begin{cases}AX_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}X_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ X_{i}^{(k)}=\Pi_{i-1}^{(k-1)},&\textrm{at }x_{i-1},i\neq 1,\\ X_{i}^{(k)}=\Pi_{i}^{(k-1)},&\textrm{at }x_{i},i\neq N,\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}A\Xi_{i}^{(k)}=\kappa_{i}\frac{d^{2}}{dx^{2}}\Xi_{i}^{(k)},&\textrm{in}\;(x_{i-1},x_{i}),\\ -\kappa_{i}\frac{d}{dx}\Xi_{i}^{(k)}=\kappa_{i-1}\frac{d}{dx}X_{i-1}^{(k)}+\kappa_{i}\frac{d}{dx}X_{i}^{(k)},&\textrm{at }x_{i-1},i\neq 1,\\ \kappa_{i}\frac{d}{dx}\Xi_{i}^{(k)}=\kappa_{i}\frac{d}{dx}X_{i}^{(k)}+\kappa_{i+1}\frac{d}{dx}X_{i+1}^{(k)},&\textrm{at }x_{i},i\neq N.\end{cases}\end{array}$
The formula is then used to update the interface values for $i=1,\ldots,N-1$ becomes:
(44) $\Pi_{i}^{(k)}$ $=\Pi_{i}^{(k-1)}-\begin{bmatrix}\theta_{i}I&0\\ 0&\phi_{i}I\end{bmatrix}\left(\Xi_{i}^{(k)}(x_{i})+\Xi_{i+1}^{(k)}(x_{i})\right)$
A quick calculation of the Dirichlet subproblems in (43) for $i=1,\ldots,N$ gives:
(45) $X^{(k)}_{i}=P\sinh\left((x-x_{i-1})\sqrt{\Lambda/\kappa_{i}}\right)\operatorname{cosech}\left(h_{i}\sqrt{\Lambda/\kappa_{i}}\right)P^{-1}\Pi_{i}^{(k-1)}$
$+P\sinh\left((x_{i}-x)\sqrt{\Lambda/\kappa_{i}}\right)\operatorname{cosech}\left(h_{i}\sqrt{\Lambda/\kappa_{i}}\right)P^{-1}\Pi_{i-1}^{(k-1)}.$
And using the solutions (45), Neumann subproblems in (43) give:
$\hat{\Xi}^{(k)}_{1}(x,s)$ $=P\sinh\left((x-x_{0})\sqrt{\Lambda/\kappa_{1}}\right)C^{(k-1)}_{1},$
$\hat{\Xi}^{(k)}_{i}(x,s)$ $=P\cosh\left((x-x_{i-1})\sqrt{\Lambda/\kappa_{i}}\right)C^{(k-1)}_{i}+P\cosh\left((x_{i}-x)\sqrt{\Lambda/\kappa_{i}}\right)D^{(k-1)}_{i},\mbox{for $i=2,\ldots,N-1,$}$
$\hat{\Xi}^{(k)}_{N}(x,s)$ $=P\sinh\left((x_{N}-x)\sqrt{\Lambda/\kappa_{N}}\right)D^{(k-1)}_{N}.$
To make things simpler, we define $\sigma_{i}:=\sinh\left(a_{i}\sqrt{\Lambda/\kappa_{i}}\right)$ and $\gamma_{i}:=\cosh\left(a_{i}\sqrt{\Lambda/\kappa_{i}}\right)$. Then putting the boundary values in Neumann subproblems, we have:
$C^{(k-1)}_{1}$ $=\gamma_{1}^{-1}\left(\left(\gamma_{1}\sigma_{1}^{-1}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}\gamma_{2}\sigma_{2}^{-1}\right)P^{-1}\Pi_{1}^{(k-1)}-\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}\sigma_{2}^{-1}P^{-1}\Pi_{2}^{(k-1)}\right),$
$C^{(k-1)}_{i}$ $=\sigma_{i}^{-1}\left(\left(\gamma_{i}\sigma_{i}^{-1}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\gamma_{i+1}\sigma_{i+1}^{-1}\right)P^{-1}\Pi_{i}^{(k-1)}-\sigma_{i}^{-1}P^{-1}\Pi_{i-1}^{(k-1)}-\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\sigma_{i+1}^{-1}P^{-1}\Pi_{i+1}^{(k-1)}\right),\quad i=2,\ldots,N-1,$
$D^{(k-1)}_{i}$ $=\sigma_{i}^{-1}\left(\left(\gamma_{i}\sigma_{i}^{-1}+\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\gamma_{i-1}\sigma_{i-1}^{-1}\right)P^{-1}\Pi_{i-1}^{(k-1)}-\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\sigma_{i-1}^{-1}P^{-1}\Pi_{i-2}^{(k-1)}-\sigma_{i}^{-1}P^{-1}\Pi_{i}^{(k-1)}\right),\quad i=2,\ldots,N-1,$
$D^{(k-1)}_{N}$ $=\gamma_{N}^{-1}\left(\left(\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}\gamma_{N-1}\sigma_{N-1}^{-1}+\gamma_{N}\sigma_{N}^{-1}\right)P^{-1}\Pi_{N-1}^{(k-1)}-\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}\sigma_{N-1}^{-1}P^{-1}\Pi_{N-2}^{(k-1)}\right).$
Substituting $\Xi_{i}$ values in updating equation (44) and using the identity $\gamma_{i}^{2}-\sigma_{i}^{2}=I$, we have
(46) $\Pi_{1}^{(k)}$ $=\Pi_{1}^{(k-1)}-\begin{bmatrix}\theta_{1}I&0\\ 0&\phi_{1}I\end{bmatrix}P\left(\left(2I+\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}\gamma_{1}\sigma_{1}^{-1}\gamma_{2}\sigma_{2}^{-1}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}\sigma_{1}\gamma_{1}^{-1}\gamma_{2}\sigma_{2}^{-1}\right)P^{-1}\Pi_{1}^{(k-1)}\right.$
$\hskip 10.0pt\left.\kern-1.2pt+\sigma_{2}^{-1}\left(\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\gamma_{3}\sigma_{3}^{-1}-\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}\sigma_{1}\gamma_{1}^{-1}\right)P^{-1}\Pi_{2}^{(k-1)}-\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\sigma_{2}^{-1}\sigma_{3}^{-1}P^{-1}\Pi_{3}^{(k-1)}\right),$
(47) $\Pi_{i}^{(k)}$ $=\Pi_{i}^{(k-1)}-\begin{bmatrix}\theta_{i}I&0\\ 0&\phi_{i}I\end{bmatrix}P\left(\left(2I+\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\gamma_{i}\sigma_{i}^{-1}\gamma_{i+1}\sigma_{i+1}^{-1}\right)P^{-1}\Pi_{i}^{(k-1)}\right.$
$\hskip 2.0pt\left.\kern-1.2pt+\sigma_{i+1}^{-1}\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\gamma_{i+2}\sigma_{i+2}^{-1}-\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\gamma_{i}\sigma_{i}^{-1}\right)P^{-1}\Pi_{i+1}^{(k-1)}+\sigma_{i}^{-1}\left(\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\gamma_{i-1}\sigma_{i-1}^{-1}-\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}\gamma_{i+1}\sigma_{i+1}^{-1}\right)P^{-1}\Pi_{i-1}^{(k-1)}\right.$
$\hskip 10.0pt\left.\kern-1.2pt-\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\sigma_{i+1}^{-1}\sigma_{i+2}^{-1}P^{-1}\Pi_{i+2}^{(k-1)}-\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\sigma_{i}^{-1}\sigma_{i-1}^{-1}P^{-1}\Pi_{i-2}^{(k-1)}\right),\quad i=2,\ldots,N-2,$
(48) $\Pi_{N-1}^{(k)}$ $=\Pi_{N-1}^{(k-1)}-\begin{bmatrix}\theta_{N-1}I&0\\ 0&\phi_{N-1}I\end{bmatrix}P\left(\left(2I+\sqrt{\frac{\kappa_{N}}{\kappa_{N-1}}}\gamma_{N-1}\sigma_{N-1}^{-1}\gamma_{N}\sigma_{N}^{-1}+\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}\sigma_{N}\gamma_{N}^{-1}\gamma_{N-1}\sigma_{N-1}^{-1}\right)P^{-1}\Pi_{N-1}^{(k-1)}\right.$
$\hskip 5.0pt\left.\kern-1.2pt+\sigma_{N-1}^{-1}\left(\sqrt{\frac{\kappa_{N-2}}{\kappa_{N-1}}}\gamma_{N-2}\sigma_{N-2}^{-1}-\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}\sigma_{N}\gamma_{N}^{-1}\right)P^{-1}\Pi_{N-2}^{(k-1)}-\sqrt{\frac{\kappa_{N-2}}{\kappa_{N-1}}}\sigma_{N-1}^{-1}\sigma_{N-2}^{-1}P^{-1}\Pi_{N-3}^{(k-1)}\right).$
Using $\theta_{i}=\bar{\theta}_{i}:=1/(2+\sqrt{\kappa_{i}/\kappa_{i+1}}+\sqrt{\kappa_{i+1}/\kappa_{i}})$ and $\phi_{i}=\bar{\phi}_{i}:=1/(2+\sqrt{\kappa_{i}/\kappa_{i+1}}+\sqrt{\kappa_{i+1}/\kappa_{i}})$for super-linear convergence in (46)-(48), we have:
(49) $\Pi_{1}^{(k)}$ $=-\bar{\theta_{1}}P\left(\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}(\gamma_{1}\sigma_{1}^{-1}\gamma_{2}\sigma_{2}^{-1}-I)+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}(\sigma_{1}\gamma_{1}^{-1}\gamma_{2}\sigma_{2}^{-1}-I)\right)P^{-1}\Pi_{1}^{(k-1)}\right.$
$\hskip 10.0pt\left.\kern-1.2pt+\sigma_{2}^{-1}\left(\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\gamma_{3}\sigma_{3}^{-1}-\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}\sigma_{1}\gamma_{1}^{-1}\right)P^{-1}\Pi_{2}^{(k-1)}-\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\sigma_{2}^{-1}\sigma_{3}^{-1}P^{-1}\Pi_{3}^{(k-1)}\right),$
(50) $\Pi_{i}^{(k)}$ $=-\bar{\theta_{i}}P\left(\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\gamma_{i}\sigma_{i}^{-1}\gamma_{i+1}\sigma_{i+1}^{-1}-I\right)P^{-1}\Pi_{i}^{(k-1)}\right.$
$\hskip 2.0pt\left.\kern-1.2pt+\sigma_{i+1}^{-1}\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\gamma_{i+2}\sigma_{i+2}^{-1}-\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\gamma_{i}\sigma_{i}^{-1}\right)P^{-1}\Pi_{i+1}^{(k-1)}+\sigma_{i}^{-1}\left(\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\gamma_{i-1}\sigma_{i-1}^{-1}-\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}\gamma_{i+1}\sigma_{i+1}^{-1}\right)P^{-1}\Pi_{i-1}^{(k-1)}\right.$
$\hskip 10.0pt\left.\kern-1.2pt-\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\sigma_{i+1}^{-1}\sigma_{i+2}^{-1}P^{-1}\Pi_{i+2}^{(k-1)}-\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\sigma_{i}^{-1}\sigma_{i-1}^{-1}P^{-1}\Pi_{i-2}^{(k-1)}\right),\quad i=2,\ldots,N-2,$
(51) $\Pi_{N-1}^{(k)}$ $=-\bar{\theta_{N-1}}P\left(\left(\sqrt{\frac{\kappa_{N}}{\kappa_{N-1}}}(\gamma_{N-1}\sigma_{N-1}^{-1}\gamma_{N}\sigma_{N}^{-1}-I)+\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}(\sigma_{N}\gamma_{N}^{-1}\gamma_{N-1}\sigma_{N-1}^{-1}-I)\right)P^{-1}\Pi_{N-1}^{(k-1)}\right.$
$\hskip 5.0pt\left.\kern-1.2pt+\sigma_{N-1}^{-1}\left(\sqrt{\frac{\kappa_{N-2}}{\kappa_{N-1}}}\gamma_{N-2}\sigma_{N-2}^{-1}-\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}\sigma_{N}\gamma_{N}^{-1}\right)P^{-1}\Pi_{N-2}^{(k-1)}-\sqrt{\frac{\kappa_{N-2}}{\kappa_{N-1}}}\sigma_{N-1}^{-1}\sigma_{N-2}^{-1}P^{-1}\Pi_{N-3}^{(k-1)}\right).$
Lemma 7.1.
For $l_{0},l_{1},l_{2}>0$, and $z\in\mathbb{C}$ such that $Re(z)\geq x_{0}>0$ we have the following estimates:
  1. (i)
    $\left|\frac{\cosh(l_{1}z)\cosh(l_{2}z)}{\sinh(l_{1}z)\sinh(l_{2}z)}-1\right|\leq\coth(l_{1}x_{0})\coth(l_{2}x_{0})-1$
  2. (ii)
    $\left|\frac{\sinh(l_{1}z)\cosh(l_{2}z)}{\sinh(l_{2}z)\cosh(l_{1}z)}-1\right|\leq\coth(l_{1}x_{0})\coth(l_{2}x_{0})-1$
  3. (iii)
    $\left|\frac{1}{\sinh(l_{1}z)\sinh(l_{2}z)}\right|\leq\operatorname{cosech}(l_{1}x_{0})\operatorname{cosech}(l_{2}x_{0})$
  4. (iv)
    $\left|\frac{1}{\sinh(l_{0}z)}\left(\rho_{1}\frac{\cosh(l_{1}z)}{\sinh(l_{1}z)}-\rho_{2}\frac{\cosh(l_{2}z)}{\sinh(l_{2}z)}\right)\right|\leq\operatorname{cosech}(l_{0}x_{0})\left(\rho_{1}\coth(l_{1}x_{0})+\rho_{2}\coth(l_{2}x_{0})\right)$
  5. (v)
    $\left|\frac{1}{\sinh(l_{0}z)}\left(\rho_{1}\frac{\cosh(l_{1}z)}{\sinh(l_{1}z)}-\rho_{2}\frac{\sinh(l_{2}z)}{\cosh(l_{2}z)}\right)\right|\leq\operatorname{cosech}(l_{0}x_{0})\left(\rho_{1}\coth(l_{1}x_{0})+\rho_{2}\coth(l_{2}x_{0})\right)$
Proof 7.2.
The procedure to prove all five parts is the same, with the only difference being the combination of exponentials. Thus, we will present the proof up to certain standard levels for each case, after which they will follow the same procedure.
  1. (i)
    Rewrite the hyperbolic terms in exponential we obtain:
    $\left|\frac{\cosh(l_{1}z)\cosh(l_{2}z)}{\sinh(l_{1}z)\sinh(l_{2}z)}-1\right|$
    $=\left|\frac{\cosh((l_{1}-l_{2})z)}{\sinh(l_{1}z)\sinh(l_{2}z)}\right|$
    $\leq 2\left|\frac{\exp(-2l_{1}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|+2\left|\frac{\exp(-2l_{2}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|$
    Now for the first part, we get by part (i) of the Lemma LABEL:SimpleLaplaceLemma
    $\left|\frac{\exp(-2l_{1}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|$ $\leq\left|\frac{\exp(-l_{1}z)}{(1-\exp(-2l_{1}z))}\right|\left|\frac{\exp(-l_{1}z)}{(1-\exp(-2l_{2}z))}\right|$
    $\leq\frac{\exp(-l_{1}x_{0})}{(1-\exp(-2l_{1}x_{0}))}\frac{\exp(-l_{1}x_{0})}{(1-\exp(-2l_{2}x_{0}))}$
    $=\frac{\exp(-2l_{1}x_{0})}{(1-\exp(-2l_{1}x_{0}))(1-\exp(-2l_{2}x_{0}))}$
    similarly,
    $\left|\frac{\exp(-2l_{2}z)}{(1-\exp(-2l_{1}z))(1-\exp(-2l_{2}z))}\right|\leq\frac{\exp(-2l_{2}x_{0})}{(1-\exp(-2l_{1}x_{0}))(1-\exp(-2l_{2}x_{0}))}$
    Therefore
    $\left|\frac{\cosh(l_{1}z)\cosh(l_{2}z)}{\sinh(l_{1}z)\sinh(l_{2}z)}-1\right|\leq\frac{\cosh(l_{1}x_{0})\cosh(l_{2}x_{0})}{\sinh(l_{1}x_{0})\sinh(l_{2}x_{0})}-1.$
    Similar simplification is possible for the second part $I_{2}$. Finally, we have our result using Lemma LABEL:estimate_1 .
    Proofs of parts (ii) and (iii) are similar to part (i), hence omitted.
  2. (iv)
    Using the triangular inequality, we have:
    $\left|\frac{1}{\sinh(l_{0}z)}\left(\rho_{1}\frac{\cosh(l_{1}z)}{\sinh(l_{1}z)}-\rho_{2}\frac{\cosh(l_{2}z)}{\sinh(l_{2}z)}\right)\right|\leq\rho_{1}\left|\frac{\cosh(l_{1}z)}{\sinh(l_{0}z)\sinh(l_{1}z)}\right|+\rho_{2}\left|\frac{\cosh(l_{2}z)}{\sinh(l_{0}z)\sinh(l_{2}z)}\right|$
    Now rewrite the hyperbolic terms in exponential for $I_{3}$ give:
    $\rho_{1}\left|\frac{\cosh(l_{1}z)}{\sinh(l_{0}z)\sinh(l_{1}z)}\right|$ $=2\rho_{1}\left|\frac{\exp(-l_{0}z)+\exp(-(l_{0}+2l_{1})z)}{(1-\exp(-2l_{0}z))(1-\exp(-2l_{1}z))}\right|$
    $\leq 2\rho_{1}\frac{\exp(-l_{0}x_{0})+\exp(-(l_{0}+2l_{1})x_{0})}{(1-\exp(-2l_{0}x_{0}))(1-\exp(-2l_{1}x_{0}))}$
    $\leq\rho_{1}\frac{\cosh(l_{1}x_{0})}{\sinh(l_{0}x_{0})\sinh(l_{1}x_{0})}$
    Similarly
    $\rho_{2}\left|\frac{\cosh(l_{2}z)}{\sinh(l_{0}z)\sinh(l_{2}z)}\right|\leq\rho_{2}\frac{\cosh(l_{2}x_{0})}{\sinh(l_{0}x_{0})\sinh(l_{2}x_{0})}$
    Therefore
    $\left|\frac{1}{\sinh(l_{0}z)}\left(\rho_{1}\frac{\cosh(l_{1}z)}{\sinh(l_{1}z)}-\rho_{2}\frac{\cosh(l_{2}z)}{\sinh(l_{2}z)}\right)\right|\leq\frac{1}{\sinh(l_{0}x_{0})}\left(\rho_{1}\frac{\cosh(l_{1}x_{0})}{\sinh(l_{1}x_{0})}+\rho_{2}\frac{\cosh(l_{2}x_{0})}{\sinh(l_{2}x_{0})}\right)$
    similar for $I_{4}$. Finally, using Lemma LABEL:estimate_1 gives the result.
  3. (v)
    The proof is similar to part (iv) and thus omitted.
Now we are prepared to present the following convergence theorems:
Theorem 7.3 (Convergence of NNWR for sub-diffusion and diffusion-wave).
For $\theta_{i}=\bar{\theta}_{i}=1/(2+\sqrt{\kappa_{i}/\kappa_{i+1}}+\sqrt{\kappa_{i+1}/\kappa_{i}})$, the NNWR algorithm for the fractional order $(0<\nu\leq 1)$ in (LABEL:NNWRD)-(11) for time window $T$, follow the following bound:
$\max_{1\leq i\leq N-1}||\Pi_{i}^{(k)}||_{\infty}\leq\bar{d}^{k}\operatorname{cosech}^{2k}\left(\frac{h}{2}\lambda\right)\|P\|_{\infty}\|P^{-1}\|_{\infty}\max_{1\leq j\leq N-1}||\Pi_{i}^{(0)}||_{\infty}.$
Proof 7.4.
Define $h_{i}:=a_{i}/\sqrt{\kappa_{i}}$. Set $\hat{\upsilon}^{(k)}_{i}(s):=\sigma^{k}\hat{\omega}^{(k)}_{i}(s)$ to reduce the equations (LABEL:NN4)-(LABEL:NN6) to the form
(52) $P^{-1}\Pi_{1}^{(k)}$ $=-\bar{\theta}_{1}\left(T_{1,1}P^{-1}\Pi_{1}^{(k-1)}+T_{1,2}P^{-1}\Pi_{2}^{(k-1)}-T_{1,3}P^{-1}\Pi_{3}^{(k-1)}\right),$
$P^{-1}\Pi_{i}^{(k)}$ $=-\bar{\theta}_{i}\left(T_{i,i}P^{-1}\Pi_{i}^{(k-1)}+T_{i,i+1}P^{-1}\Pi_{i+1}^{(k-1)}+\hat{t}_{i,i-1}P^{-1}\Pi_{i-1}^{(k-1)}-T_{i,i+2}P^{-1}\Pi_{i+2}^{(k-1)}-T_{i,i-2}P^{-1}\Pi_{i-2}^{(k-1)}\right),$
$P^{-1}\Pi_{N-1}^{(k)}$ $=-\bar{\theta}_{N-1}\left(T_{N-1,N-1}P^{-1}\Pi_{N-1}^{(k-1)}+T_{N-1,N-2}P^{-1}\Pi_{N-2}^{(k-1)}-T_{N-1,N-3}P^{-1}\Pi_{N-3}^{(k-1)}\right),$
where the weights are given by
(53) $T_{1,1}=\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}(\gamma_{1}\sigma_{1}^{-1}\gamma_{2}\sigma_{2}^{-1}-I)+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}(\sigma_{1}\gamma_{1}^{-1}\gamma_{2}\sigma_{2}^{-1}-I)\right),$
$T_{1,2}=\sigma_{2}^{-1}\left(\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\gamma_{3}\sigma_{3}^{-1}-\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}\sigma_{1}\gamma_{1}^{-1}\right),\,T_{1,3}=\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\sigma_{2}^{-1}\sigma_{3}^{-1},$
for $i=2,\cdots,N-2$,
$T_{i,i}=\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\gamma_{i}\sigma_{i}^{-1}\gamma_{i+1}\sigma_{i+1}^{-1}-I\right),\,T_{i,i+1}=\sigma_{i+1}^{-1}\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\gamma_{i+2}\sigma_{i+2}^{-1}-\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\gamma_{i}\sigma_{i}^{-1}\right),$
$T_{i,i-1}=\sigma_{i}^{-1}\left(\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\gamma_{i-1}\sigma_{i-1}^{-1}-\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}\gamma_{i+1}\sigma_{i+1}^{-1}\right),\,T_{i,i+2}=\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\sigma_{i+1}^{-1}\sigma_{i+2}^{-1},\,T_{i,i-2}=\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\sigma_{i}^{-1}\sigma_{i-1}^{-1},$
and
$T_{N-1,N-1}=\left(\sqrt{\frac{\kappa_{N}}{\kappa_{N-1}}}(\gamma_{N-1}\sigma_{N-1}^{-1}\gamma_{N}\sigma_{N}^{-1}-I)+\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}(\sigma_{N}\gamma_{N}^{-1}\gamma_{N-1}\sigma_{N-1}^{-1}-I)\right),$
$T_{N-1,N-2}=\sigma_{N-1}^{-1}\left(\sqrt{\frac{\kappa_{N-2}}{\kappa_{N-1}}}\gamma_{N-2}\sigma_{N-2}^{-1}-\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}\sigma_{N}\gamma_{N}^{-1}\right),\,T_{N-1,N-3}=\sqrt{\frac{\kappa_{N-2}}{\kappa_{N-1}}}\sigma_{N-1}^{-1}\sigma_{N-2}^{-1}.$
These weights $\hat{t}_{i,j}$ in (LABEL:NNWR3) are in the same form as in Lemma LABEL:lam_7. Therefore, we have the following estimates:
for $i=1,2,\cdots,N-1$,
$||T_{i,i}(.)||_{\infty}$ $\leq\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\coth(h_{i}\lambda)\coth(h_{i+1}\lambda)-1\right)=:w_{i,i},$
for $i=1,\cdots,N-2,$
$||T_{i,i+1}(.)||_{\infty}$ $\leq\operatorname{cosech}(h_{i+1}\lambda)\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\coth(h_{i+2}\lambda)+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\coth(h_{i}\lambda)\right)=:w_{i,i+1},$
$||T_{i,i+2}(.)||_{\infty}$ $\leq\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\operatorname{cosech}(h_{i+1}\lambda)\operatorname{cosech}(h_{i+2}\lambda)=:w_{i,i+2},$
for $i=2,\cdots,N-1,$
$||T_{i,i-1}(.)||_{\infty}$ $\leq\operatorname{cosech}(h_{i}\lambda)\left(\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\coth(h_{i-1}\lambda)+\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}\coth(h_{i+1}\lambda)\right)=:w_{i,i-1},$
$||T_{i,i-2}(.)||_{\infty}$ $\leq\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\operatorname{cosech}(h_{i-1}\lambda)\operatorname{cosech}(h_{i}\lambda)=:w_{i,i-2}.$
After having the inverse Laplace transform in (LABEL:NN7) and using part (ii) of Lemma LABEL:SimpleLaplaceLemma we obtain
$||P^{-1}\Pi_{1}^{(k)}||_{\infty}$ $\leq c_{1}\bar{\theta}_{1}\max_{1\leq j\leq N-1}||P^{-1}\Pi_{j}^{(k-1)}||_{\infty},$
(54) $||P^{-1}\Pi_{i}^{(k)}||_{\infty}$ $\leq c_{i}\bar{\theta}_{i}\max_{1\leq j\leq N-1}||P^{-1}\Pi_{j}^{(k-1)}||_{\infty},$
$||P^{-1}\Pi_{N-1}^{(k)}||_{\infty}$ $\leq c_{N-1}\bar{\theta}_{N-1}\max_{1\leq j\leq N-1}||P^{-1}\Pi_{j}^{(k-1)}||_{\infty}.$
where
$c_{1}$ $=w_{1,1}+w_{1,1+1}+w_{1,1+2},$
$c_{i}$ $=w_{i,i}+w_{i,i-1}+w_{i,i+1}+w_{i,i-2}+w_{i,i+2},\quad i=2,\cdots,N-2,$
$c_{N-1}$ $=w_{N-1,N-1}+w_{N-1,N-2}+w_{N-1,N-3}.$
Define
(55) $\bar{c}:=\max_{1\leq i\leq N-1}\bar{\theta}_{i}c_{i}.$
Then from equations (LABEL:NNWR4), we have
(56) $||P^{-1}\Pi_{i}^{(k)}||_{\infty}\leq\bar{c}\max_{1\leq j\leq N-1}||P^{-1}\Pi_{i}^{(k-1)}||_{\infty},$
which further yields
$\max_{1\leq i\leq N-1}||\Pi_{i}^{(k)}||_{\infty}\leq\bar{c}^{k}\|P\|_{\infty}\|P^{-1}\|_{\infty}\max_{1\leq j\leq N-1}||\Pi_{i}^{(0)}||_{\infty},$
we can find the bound in terms of minimum subdomain length. For that:
$w_{i,i}$ $\leq\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\coth(h_{i}\lambda)\coth(h_{i+1}\lambda)-1\right)$
$\leq\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\left(\coth^{2}(h\lambda)-1\right)$
$=\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\operatorname{cosech}^{2}(h\lambda)$
Similarly,
$w_{i,i+1}$ $\leq\left(\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}\right)\operatorname{cosech}(h\lambda)\coth(h\lambda)$
$w_{i,i+2}$ $\leq\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}\operatorname{cosech}^{2}(h\lambda)$
$w_{i,i-1}$ $\leq\left(\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}+\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}\right)\operatorname{cosech}(h\lambda)\cot(h\lambda)$
$w_{i,i-2}$ $\leq\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\operatorname{cosech}^{2}(h\lambda).$
Therefore
$c_{1}$ $\leq\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}+\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\right)\operatorname{cosech}(h\lambda)(\operatorname{cosech}(h\lambda)+\coth(h\lambda))$
$=\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}+\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\right)\frac{2\cosh^{2}(\frac{h}{2}\lambda)}{\sinh^{2}(h\lambda)}$
$=\left(\sqrt{\frac{\kappa_{1}}{\kappa_{2}}}+\sqrt{\frac{\kappa_{2}}{\kappa_{1}}}+\sqrt{\frac{\kappa_{3}}{\kappa_{2}}}\right)\frac{\operatorname{cosech}^{2}(\frac{h}{2}\lambda)}{2}$
$=:d_{1}\operatorname{cosech}^{2}\left(\frac{h}{2}\lambda\right)$
Similarly
$c_{i}$ $=\left(\sqrt{\frac{\kappa_{i}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i+1}}{\kappa_{i}}}+\sqrt{\frac{\kappa_{i+2}}{\kappa_{i+1}}}+\sqrt{\frac{\kappa_{i-1}}{\kappa_{i}}}\right)\frac{\operatorname{cosech}^{2}(\frac{h}{2}\lambda)}{2}$
$=:d_{i}\operatorname{cosech}^{2}\left(\frac{h}{2}\lambda\right).$
$c_{N-1}$ $=\left(\sqrt{\frac{\kappa_{N-1}}{\kappa_{N}}}+\sqrt{\frac{\kappa_{N}}{\kappa_{N-1}}}+\sqrt{\frac{\kappa_{N-2}}{\kappa_{N-1}}}\right)\frac{\operatorname{cosech}^{2}(\frac{h}{2}\lambda)}{2}$
$=:d_{N-1}\operatorname{cosech}^{2}\left(\frac{h}{2}\lambda\right).$
Therefore:
$\max_{1\leq i\leq N-1}||\Pi_{i}^{(k)}||_{\infty}\leq\bar{d}^{k}\operatorname{cosech}^{2k}\left(\frac{h}{2}\lambda\right)\|P\|_{\infty}\|P^{-1}\|_{\infty}\max_{1\leq j\leq N-1}||\Pi_{i}^{(0)}||_{\infty}.$
where $\bar{d}:=\max_{1\leq i\leq N-1}\bar{\theta}_{i}d_{i}$.